A Complete Bibliography of Publications in Fisheries Research (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

18 September 2024 Version 1.05

Title word cross-reference

$10.90 [Ken85]. $11.00 [Lan85]. $11.50 [Fur85b]. $12.50 [Bor86, Bro83]. $15.00 [Haw82]. $16.00 [Fur85a, Thr85]. $16.70 [Mit85]. $17.50 [Mit82b, Som89]. $18 [Paw89]. $19.00 [Wis84]. $19.50 [Thr85]. $2.50 [Mit82a]. $24.00 [Mag86]. $25.00 [Bor86, Kes84a]. $28.00 [Da 85a, Da 85b, Nau85, Pil86]. $31.90/£ [Fra86]. $32.00 [Tho86b]. $32.50 [Gjø84a]. $32.95 [Mit87]. $33.00 [Thr85]. $35.00 [Joh85]. $36.00 [Som89]. $37.00 [Thr85]. $37.70 [Wym86]. $4.50 [Kes84d]. $42.30 [Lon85]. $43.00 [Mit82b]. $47.50 [Cha82]. $49.50 [Wei85]. $5.00 [Eve84]. $50.90 [Lew85]. $55.00 [Smi83]. $6.55 [Ken85]. $600.00 [Kes82b]. $63.50 [Daa87a]. $66.75./Dfl [Lee88]. $69.95 [Wid86]. $7.00 [Fur85b]. $78.00 [Kes85a]. $8.00 [You86]. $8.50 [All89]. $9.00 [Fur85a]. $9.50 [Paw89]. $A [Thr84]. × [Cus82a, Dah82a, Dah82b, Lis82, Smi82, Suo82]. x [Mit85, Pri84b].

0 [Ano86c, Bor86, Car83, Car86b, Cus82a, Dah82a, Dah82b, De 86, Fra86, Fre89, Hop89, Kar84, Kes82a, Kes82c, Kes82d, Kes83a, Kes83c, Kes85b, Lew85, Lin87, Lis82, Loc88, Lon85, Mac83, Mit82b, Mit88, Moh83, Nak85, Pee84, Pol83, Pri89, Ruc86, Sae82, Smi82, Som87, Suo82, Tho86b, Tho88, Tho89b, Thr84, UB89b, Wid86, All83, Hee89, Joh85, Kes83b, Lee88, Mit84, Mit87, Pri84b, Tho89a, Wei85, Wis84]. 0-0471-91114-3 [Fre89]. 0-216-91017-X [Pol83]. 0-295-95768-9 [All83]. 0-295-95982-7 [Mit84]. 0-304-30315-1 [Kes82c]. 0-444-42040-1 [Pri84b]. 0-444-42736-8 [Lee88]. 0-471-89236-X [Som87]. 0-471-90027-3 [Fra86]. 0-520-03929-7 [Kes83c]. 0-521-24303-3 [Wei85]. 0-7099-1761-9 [Ano86c]. 0-7099-1837-2 [Lin87]. 0-7099-3344-4 [Mit88]. 0-7099-4786-0 [Pri89]. 0-7992-0465-X [Mac83]. 0-8014-1334-6 [Kes83b]. 0-8018-3253-5 [Mit87]. 0-8133-7145-7 [Hee89]. 0-85238-061-5 [Dah82a]. 0-85238-108-5 [Kes82a]. 0-85238-110-7 [Suo82]. 0-85238-111-5 [Dah82b]. 0-85238-112-3 [Kes82d]. 0-85238-115-8 [Moh83]. 0-85238-116-6 [Cus82a]. 0-85238-117-4 [Smi82]. 0-85238-118-2 [Kar84]. 0-85238-119-0 [Car83]. 0-85238-120-4 [Lis82]. 0-85238-122-0 [Pee84]. 0-85238-124-7 [Nak85]. 0-85238-125-5 [Car86b]. 0-85238-128-X [Kes85b]. 0-85238-129-8 [Ruc86]. 0-85238-133-6 [Hop89]. 0-85238-141-7 [Tho88]. 0-85238-144-1 [Loc88]. 0-85664-894-9 [Gjø84a]. 0-85976-125-8 [De 86]. 0-86346-041-0 [Bor86]. 0-86346-042-9 [Bor86]. 0-86840-129-3 [Thr84]. 0-87055-341-0 [Wis84]. 0-87055-405-0 [Gjø84a]. 0-87055-438-7 [Joh85]. 0-87371-089-4 [UB89b]. 0-87371-090-8 [UB89b]. 0-87371-091-6 [UB89b]. 0-87371-092-4 [UB89b]. 0-87451-451-7 [Tho89b]. 0-88936-336-6 [Kes83a]. 0-89116-302-6 [Wid86]. 0-89158-353-X [Mit82b]. 0-89158-683-0 [Sae82]. 0-913235-00-8 [Tho86b]. 0-913235-14-8 [Lan85]. 0-939923-05-X [Tho89a]. 0.12.567960.2 [Wym84]. 0097 [Lan85]. 0097-0638 [Lan85]. 0115 [Cus82b, Fur85a, Fur85b, Ken85, Tro82, All89]. 0115-4389 [Cus82b, Tro82]. 0115-5547 [Fur85a, Fur85b, Ken85, All89]. 0116 [McI89a]. 0116-6514 [McI89a]. 0638 [Lan85]. 0738 [Kes85a]. 0738-1360 [Kes85a].

1 [Kes82c, Loc88, Pet86, Smi83, Pri84b, Som88]. 1360 [Kes85a]. 16.50/US [Bor86]. 17.50 [Pil86]. 1979 [Wym83, Wym82]. 1984 [Kir86]. 1987 [McI89a]. 198e [Kes83a]. 1er [Kes84c].

2 [Kar84, Lin87, Daa87a]. 20-year [GM88]. 25th [Kir83]. 26.60/$44.00 [Wym84]. 29th [Kir83]. 2nd [Kar84, Kes83b, Wym83].

3 [Da 85a, Fin86, Fra86, Fre89, Gar84, Kes82d, Kir86, Lew85, Lon85, McI89b, Mit85, Nae86, Paw89, Pri86, Smi83, Som86, Thr84, Wal89, Wym86, Kes84b, Wei85]. 3-540-11583-8 [Kes84b]. 3-540-11654-0 [Lew85]. 3-540-12013-0 [Lon85]. 3-540-12698-8 [Mit85]. 3-540-13707-6 [Pri86]. 3-540-15028-5 [Som86]. 3-540-15982-7 [Kir86]. 3-540-16851-6 [McI89b]. 3-540-90590-1 [Smi83]. 3-540-90935-4 [Wym86]. 3-540-96084-8 [Nae86]. 3-8055-4253-4 [Wal89]. 3rd [Car86b].

4 [Da 85b, Lis82, Mit88, Smi82, UB89b, Wal89, Wym86]. 415pp [Kir83].

43/US$15.10 [Fin86]. 4389 [Cus82b, Tro82]. 49-004 [Kes82b]. 4DY [Hee89].

5 [Car86b, Dah82a, Dah82b, Kes82a, Som86, Mit87]. 5547 [Fur85a, Fur85b, Ken85, All89]. 5AA [De 86].

6 [Bai86, Cus82a, Daa87b, Hop89, Kes83a, McI89b, Pri86, Wid86, Kes83b, Som89]. 60p [Pee84]. 65.00/U.S.$34.00 [Wym83]. 6514 [McI89a].

7 [Kes83c, Kir86, Nak85, Suo82, Tho88, Tho89b, Hee89, Joh85, Mit84]. 74271 [Nau85].

8 [Cha82, De 86, Mit85, Moh83, Nae86, Ruc86, Tho86b, Gjø84b, Kes84b, Lee88]. 8.50/US [Bor86]. 82 [Nau85]. 82-74271 [Nau85]. 85/$34/£ [Som88].

9 [All82a, Ano86c, Bor86, Car86a, All83, Mag86]. 90 [All82a, Cha82, Da 85a, Da 85b, Gjø84b, Mag86, Som88, Wym83]. 90-6191-409-4 [Da 85b]. 90-6191-421-3 [Da 85a]. 90-6191-432-9 [Mag86]. 90-6191-484-1 [Som88]. 90-6193-701-9 [All82a]. 90-6193-755-8 [Cha82]. 90-6193-821-X [Wym83]. 90-6193-822-8 [Gjø84b]. 90.00/US [Cha82]. 92 [Daa87b, Daa87a]. 92-3-002215-2 [Daa87a]. 92-3-002308-6 [Daa87b]. 971 [Bai86, Gar84, Paw89, Som89]. 971-02-0003-8 [Haw82]. 971-0400-22-3 [Gar84]. 971-10-2200-1 [Thr85]. 971-10-2201-X [Thr85]. 971-10-2203-6 [Bai86]. 971-10-2226-5 [All89]. 971-10-2230-3 [Paw89]. 971-10-2234-6 [Som89]. 978 [Fin86]. 978-0-387-96062-3 [Fin86].

A.D. [Som88]. abalone [MS89]. abundance [AH86, ANM87, BWL87, BC89, Dic89a, God89, MWS83, UB89a, Van82]. Academia [Kes82b]. Academic [Wym84]. acclimation [BG85]. account [Tan85]. Acheson [Tho89b]. Acid [Nau85]. Acidic [Nau85, BG85]. acoustic [BEKV82, Foo82, God89, UB89a]. across [Mat88]. activity [LS89, TTA82, WB82]. adaptation [SL87]. additional [Tho89a]. adjacent [RP82]. Adjusting [Jen82, PM86]. adult [App88]. advances [Lin87, Kes82a]. Aegean [TT89]. affecting [BC83]. Africa [Mac83, TBH85, TB88]. after [GM88]. Age [App88, Cro89, DG86, MD87, MM85, Pap84, PB86, TT89, Apr87, BC88, EO89, HT86, SVQ85, WHW88]. age-class [BC88]. age-distributions [HT86]. ageing [LG87]. aggregations [Foo82]. Agricultural [Suo82]. Agriculture [Fra86]. air-bubble [PCS^{+}85]. airmail [All89, Fur85a, Fur85b, Ken85, Paw89, Som89, Thr85]. Al [NNAS89]. Al-Hammar [NNAS89]. Alaska [You86, ANM87, Fag88]. Alaskan [Lan82]. alba [GM88, SGB87]. Aldershot [Hee89]. Alexander [Hee89]. alkaline [BG85]. along [MM85, TL83]. alter [FS89]. Amazon [Cha82]. America [Gjø84a, Nau85]. American [Lan85, Nau85, Tho86b, BD87, DB89]. americana [SRRU82]. americanus [BD87, DB89]. among [Jør85]. Amsterdam [Lee88, Pri84b]. analyses [MYH86]. Analysis [GM82, SY84, Som87]. analytic [Fag88]. analyzing [RLH88]. anchoveta [Som89]. anchovy [EO89]. ancient [AP86]. Anderson [Mit87]. Angeles [Kes83c]. angler [Cro87, Cro89]. angler-fish [Cro87, Cro89]. angles [PW85, PW86]. Anguilla [Apr87, Mor86, PB86]. anisakid [CC85a]. Announcement [Ano83a, Ano84a, Ano86a, Ano88a, Ano88b, Ano89a]. Announcements [Ano84b]. annual [McI89a]. Ano [Kes82b]. Anthony [All83]. Antonio [Fur85a]. Apdo [Kes82b]. aperture [LT89]. appearance [TTA82]. appendixes [Fre89]. application [Pak89, LSM85]. applications [Lew85]. approach [Fag88]. approx [Fra86, Gjø84b, Kes84b, Lew85, Lon85, Mit85, Smi83, Wym86]. April [Ano84d]. Aquaculture [Pri84b, Ruc86]. Aquatic [All89, Bai86, Haw82, Ken85, Lew85, Paw89, Som89, Wid86]. Arabian [AH86, JA89]. Arctic [GM87, NR87, SL87, Ull87]. areas [Atk89, CCS86, Dav86, Bor86]. Arno [TTA82]. arrangement [Dah82b]. Arsenic [JA89]. artificial [BCE86]. Ashland [Kir86]. Asian [McI89a, McI89a]. aspect [PW85, PW86]. Aspects [KFM87, Tho86a, TL83, AH84, PW85]. assemblage [VWN87]. assess [PDS86]. Assessment [AP86, DP89, Fra86, GS87, MS87, SRRU82, ZW88, Fur85a]. association [Joh84]. Atlantic [Daa87a, Daa87b, Atk89, Blo87, BC89, CCS86, GR82, GS87, HT86, HH88]. Atlas [Bro83]. attack [PW85, PW86]. attractant [DB89]. August [Ano87f, Nau85]. Australia [Gar84, Dav86, GM88, LSC85, MS89, SGB87, SBPR89, WHW88, WD87, WH89]. Australian [Edw83]. australiensis [WD87]. Author [Ano82a, Ano85a, Ano86b, Ano87a, Ano89b]. Authorindex [Ano88c]. autolysis [MAN86]. availability [Dic89a]. Available [Joh84, Som89, GSC89, Paw89]. AVI [Gjø84a, Joh85, Wis84].

B [Lan85, Loc88, Nak85, Pol83, Wis84, Wym82]. B.V [All82a]. Bacterial [Eve84]. Bailey [Fur85b, Loc88, Paw89]. Bait [Wis84]. baited [FH83]. balance [Bla85]. Balkema [Da 85a, Da 85b, Mag86, Pil86, Som88]. Balon [All82a]. Baltic [Ane89, KFM87]. Baltimore [Mit87]. Bank [BC89]. banks [Ber88b]. Banon [All89]. Baram [Wat82]. Barbus [NNAS89]. Bardach [Haw82]. Barents [BJD87, EW87, NR87]. Barlow [Moh83]. barnacle [Ber88a]. Barnegat [Wym86]. barramundi [Dav86]. barrier [Goe82, PCS^{+}85]. barriers [Ste82]. based [Orł89]. Basel [Wal89]. basic [Fra86]. basis [Zal85]. basking [Pri84a]. Bass [WHW88, BYNB87, Lan85]. Bate [TL83]. Bauchot [Daa87a, Daa87b]. Bay [Fur85a, Fur85b, GM88, Nae86, SGB87, WH89, Wym86, Tan85]. bays [DGMO88, GOMH85]. be [Tho89a]. beam [GE86, RLH88, War89]. Beckenham [Ano86c, Lin87, Pri89]. behavioral [Wym83]. behaviour [FH83, Haw82, NC89]. Bell [Mit82b]. Ben [Dah82b]. Ben-Yami [Dah82b]. benthic [BCE86, TL83]. Bering [You86, BWL87]. Berkeley [Kes83c]. Berlin [Fin86, Kes84b, Kir86, Lew85, Lon85, Mit85, Nae86, Pri86, Smi83, Som86, Wym86]. Bertalanffy [DGMO88]. Bethesda [Lan85, Nau85, Tho86b]. between [PC89, RLH88, Zal85]. bias [EW87, JK87]. biennial [Wym83]. biological [Sch87]. Biology [Fur85a, Gjø84b, Pol83, Wym83, AH84, GS87, Pap84, Pri89, Tho86a, Wal89, Fre89, Pol83]. biomass [GSC89, Lan82, ML88, Orł89, Zal85]. biosynthetic [Fre89]. black [DGMO88, TB88, EO89]. Blackie [Pol83]. Blacksburg [Tho89a]. Blake [Wei85]. Bleeker [DP89]. Bloch [Dav86]. blue [Tan85]. bluefin [Tho86a]. Bo [Bor86, Lee88]. board [PW85, PW86, Ano82b, Ano83b, Ano85b, Ano86d, Ano87b, Ano87c, Ano89f, Ano89g]. Boats [Pee84]. Bombay [Kes84a]. Bone [Pol83]. bonga [Mos88]. Bonus [Kes83a]. Book [All82a, All83, All89, Ano86c, Bai86, Bor86, Bro83, Car83, Car86a, Car86b, Cha82, Cus82a, Cus82b, Da 85a, Da 85b, Daa87a, Daa87b, Dah82a, Dah82b, De 86, Eve84, Fin86, Fla82, Fra86, Fre89, Fur85a, Fur85b, Gar84, Gjø84a, Gjø84b, Haw82, Hee89, Hop89, Joh84, Joh85, Kar84, Ken85, Kes82a, Kes82b, Kes82c, Kes82d, Kes83a, Kes83b, Kes83c, Kes84a, Kes84b, Kes84c, Kes84d, Kes85a, Kes85b, Kir83, Kir86, Lan85, Lee88, Lew85, Lin87, Lis82, Loc88, Lon85, Mac83, Mag86, McI89a, McI89b, Mit82b, Mit82a, Mit84, Mit85, Mit87, Mit88, Moh83, Nae86, Nak85, Nau85, Paw89, Pee84, Pet86, Pil86, Pol83, Pri84b, Pri86, Pri89, Ruc86, Sae82, Smi82, Smi83, Som86, Som87, Som88, Som89, Suo82, Tho86b, Tho88, Tho89b, Tho89a, Thr84]. Book [Thr85, Tro82, UB89b, Wal89, Wei85, Wid86, Wis84, Wym82, Wym83, Wym84, Wym86, You86]. Books [Car83, Car86a, Car86b, Cus82a, Dah82a, Dah82b, Hop89, Kar84, Kes82a, Kes82d, Kes85b, Lis82, Loc88, Moh83, Nak85, Paw89, Pee84, Pet86, Ruc86, Smi82, Som89, Suo82, Tho88]. borne [BEKV82]. bottom [Mat88, NR87, Ste88, VWN87]. bottom-set [Ste88]. Boulder [Mit82b, Sae82]. Bowdich [Mos88]. Bowker [Mit82b, Sae82]. Box [Bai86, Paw89, Som89]. Brandt [Car86b]. Brian [Ano86c]. Bridger [Car83]. Britain [Car86a, Cus82a, Dah82a, Dah82b, De 86, Fra86, Hee89, Kar84, Kes82a, Kes82d, Kes85b, Lis82, Mit82b, Ruc86, Sae82, Smi82, Suo82]. British [Kes85b, SLH87, Ber88a, BS87, HOE89, KFM87]. Brown [Mit84, Wis84]. bubble [PCS^{+}85, SHS87, Ste82]. bubble-curtain [SHS87]. Buckland [Kes85b]. Bukit [SY84]. Bulgaria [PJDZ85]. burning [Apr87]. BV [Cha82]. by-catch [MS85, WH89]. by-products [Moh83]. byproducts [DB89].

C [Daa87a, Daa87b, Fur85b, Haw82, McI89b, Paw89]. C. [SY84]. calcarifer [Dav86]. Calculations [Tho88]. calculators [Bai86]. California [Joh85, Kes83c, BA82, VBBM89]. Cambridge [Wei85, Wym82]. Cameroon [DP89]. Can [Bro83, FS89]. Canada [Bro83, Kes83a, SLH87]. Canadian [BC89, PDS86]. canal [HOE89]. Cancer [BCE86]. cannibalism [LSM85]. capacity [FS89, Han87]. Cape [Nae86]. Capeo [Mar85]. Capetown [Mac83]. capture [Haw82, Paw89, SH82]. carbohydrate [BG85]. Carbon [Jør85, Lee88]. Cargill [Mac83]. Caribbean [Mit85]. Carlos [Kes82b]. carp [HOE89]. Carquinez [BYNB87]. Carribean [Thr85]. Carrothers [Tho88]. Cassell [Kes82c]. Catch [Bri85, SBPR89, GSC89, Han87, Jen82, MS85, SSR89, WH89]. Catchability [War89]. catches [BS87]. catching [AO87, Car86b, YF89]. caught [Fag88, HT86, Mar85]. cell [Wal89]. cement [Dah82a]. Center [All89, Bai86, Haw82, Ken85, Paw89, Som89]. central [WB82]. Centre [Kes83a]. Cetorhinus [Pri84a]. Chairman [Nau85]. chalcogramma [ANM87, LG87]. change [Car88]. Changes [ANM87, HT86, LS89, GM88, UB89a]. Chapman [Pol83]. characteristics [Ane89, DB89, GR82, SLH87]. charge [Tho89a]. charges [Tho89a]. Charles [Joh85, Wym82]. Charrs [All82a]. Chaston [Ruc86]. chemo [DB89]. chemo-attractant [DB89]. Chemoreception [Pri84b]. Chesapeake [Tan85]. Chew [Kes84d]. Chichester [Fra86, Fre89, Som87]. Chikamasa [Pee84]. Chilean [CC85a]. China [ZW88]. Chinook [Fag88, Joh88]. clam [Kes84d]. clarki [MYH86, MYH86]. class [BC88]. clavatum [Cro87]. closed [Cri83]. cloth [All82a, Da 85a, Da 85b, Lew85, Lon85, Som88, Thr85]. Clupea [Ane89, KFM87]. CO [Mit82b, Sae82, Tho89a]. coast [Kes84d, MM85, MWS83, TL83]. Coastal [Jak87, Mit85, Wym86, Fin86, GM87, MYH86]. Cod [Dic89a, Dic89b, Nae86, EW87, FH83, GM87, God89, Jak87, Jør85, LSC85, NR87, SL87, Ull87]. cod-end [LSC85]. codend [SF87]. codends [FS87]. Coho [Joh88]. Cohort [SY84]. collected [MD87]. collectors [BCE86]. coloquio [Kes84c]. Colorado [WB82]. colour [All82a]. Columbia [All83, SLH87, Ber88a, BS87]. combinations [SHS87]. comments [Pap84]. commercial [DS89, JA89]. commercial-scale [DS89]. commercially [All89, SLH87, SJS^{+}89]. communication [Smi83]. Communications [Kes84d]. communities [Som86]. Company [Joh85, Lee88, Wis84]. Comparative [DP89, BK89, GS87]. Comparison [AH84, GW89, SVQ85, HH87, RLH88, SH82, TB88]. compensated [DGMO88]. competition [SBPR89]. Compiled [Nau85]. Composition [RP82, Bri85, EO89]. computer [Cru86]. Concentration [DB89]. concentrations [JA89]. conceptual [SJ86]. Conch [App88]. condition [BWL87, LPE83]. conditions [GM87, UB89a]. Conference [Haw82]. Connecticut [Wis84]. Connell [Kes82a, Lis82]. conservation [Blo87]. considerations [BK89]. contamination [UB89b]. Contents [Ano88d, Ano89c]. Contracaecum [Cro87]. control [BD87, Pri86]. copy [Tho89a]. Coral [Thr84, Thr85]. corals [VWN87]. Coregnous [SVQ85]. Coregonus [SN89]. Cornell [Kes83b, Nau85]. Corrigendum [Ano89d, Ano89e]. Coryphaenoides [Atk89]. cost [PDS86, WW82]. costs [BH86, WW82]. countries [Tho89a]. cover [Fin86, Mit85, Wym86]. crab [BCE86, Tan85]. crabs [SH82]. craft [Dah82a]. Crane [Kes85a]. Crayfish [Joh85, Joh85]. creel [SH82]. Cronulla [Gar84]. Croom [Ano86c, Gjø84a, Lin87, Mit88, Pri89]. Crosa [Eve84]. crowding [LPE83]. cruise [GG84]. Crustacea [TL83]. crustaceans [WD87]. Crutchfield [Mit84]. CSIRO [Gar84]. CT [Gjø84a, Joh85]. culture [Haw82]. current [BWL87]. curtain [SHS87]. curtains [SHS87, Ste82]. curves [Jen82, ST82]. cutthroat [MYH86]. CVPA [LSM85]. cycle [TL83]. cygnus [BC83]. cyprinid [NNAS89]. Cyprinidae [SY84]. cyprinids [BC88, DS87, DS89, SD89, TBH85].

D [All89, Bai86, Fre89, Gar84, Kes82b, Kes83b, Kes84b, Kes84d, Kir83, Mit82a, Nae86, Paw89, Som89, Thr84, Wym83]. D-1000 [Nae86]. D-6070 [Paw89, Som89]. Dahlem [Som86]. daily [TTA82]. damming [WR85]. Daniel [Cus82b, Fur85a]. data [BEKV82, Fag88, GSC89, RLH88]. data-analytic [Fag88]. David [Gjø84b, Lan85, Tho86b]. Davis [Joh85, Nae86, BC89]. DC [Mit82a]. Decapoda [TL83]. Decapterus [ZW88]. December [Ano86h, McI89a]. deep [GG84, WD87]. deep-water [GG84, WD87]. del [Som89]. DeLury [CC85b, Cri83]. demersal [GM82, GE86, Mac82, PM86]. demography [TBH85]. Density [RLH88, AO87, JK87, Zal85]. Departmento [Kes84c]. derbies [Fag88]. Derecho [Kes82b, Kes82b]. derived [Apr87]. Description [Sch87]. design [BS87, Hop89]. designed [PDS86]. designs [Tho88]. detailed [Fre89]. determination [App88, Apr87, BC88, PB86, SVQ85, TT89]. determine [God89]. deterministic [Han87]. Deutsche [Som89]. developing [Jør85]. Development [GE86, Kes83a, Wal89, Ken85, Som88]. Developments [Cha82, Gjø84b, Pri84b, Mac83, Loc88]. Dfl [Cha82, Da 85a, Da 85b, Daa87b, Gjø84b, Pil86, Pri84b, Som88, Wym83]. Dfl.400.00/US.$210.60 [All82a]. Dfl.68.25 [Mag86]. Dictionary [Fre89]. die-off [BYNB87]. Diel [WD87, UB89a, VBBM89]. Diemen [Dav86]. different [Atk89, GM87, Jør85]. Direct [Bor89, MS89, Tho89a]. direction [Nau85]. Director [Kes82a]. discreteness [SJ86]. diseases [Eve84]. Distributed [Hee89, Pol83]. Distribution [AH86, Blo87, BC89, Hee89, MWS83, ANM87, BC88, Jen82, NR87, NNAS89, Orł89, Van82, VBBM89]. distributions [HT86]. Distributor [Mit82b, Sae82]. Distributors [Kes84a]. district [Kes83c]. diversity [JK87]. Division [Mit82a, Nau85]. DM [Fin86, Kes84b, Lew85, Lon85, McI89b, Mit85, Pri86, Smi83, Som86, Wym86]. DM25.00 [Kir86]. DM58.00 [Nae86]. Dodecanese [PC89]. dollars [Joh84]. dolphins [Wym84]. dominated [VBBM89]. Donald [De 86]. Double [LT89]. drag [FS87, Mac82, SF87]. Draper [Wym82]. Drave [De 86]. drawdown [ST82]. drum [DGMO88, GOMH85]. Dublin [Bor86]. duration [HW85]. during [BG85, EW87, GG84]. Dwiponggo [All89, Paw89]. dynamics [BA82, Bai86, Lee88, Pak89, ZW88, Lon85].

each [Tho89a]. early [BCE86, DM87]. east [De 86, MM85]. eastern [BWL87, Daa87a, Daa87b, GG84, Mar85, TL83, You86, WR85, WHW88]. Echo [Suo82]. echoes [Foo82]. Echolocation [Wym84]. Ecological [SL87]. ecology [BJD87, Nae86, Pri89, Wym83, Gjø84a, Smi82, Wym83, Wym86]. Economic [DS88, Blo87]. Economics [Kes84a, Mit88, Mit82b, Mit87, Kes85a]. Economies [Mit84]. ecosystem [Cus82a, Som89]. ecosystems [Hee89]. Edeson [Mit85]. edible [JA89]. Edingburgh [De 86]. Edited [Kir83, Smi83, Wym83, Tho88]. Edition [Fre89, Mit82b, Sae82, Wym83]. Editor [All82a, Da 85a, Da 85b, Dah82a, Dah82b, Eve84, Hop89, Joh85, Kes82d, Kes85a, Kir86, Mac83, Nau85, Pet86, Pri84b, Som86, Som88, Thr85, UB89b, You86]. Editorial [Ano82b, Ano83b, Ano85b, Ano86d, Ano87b, Ano87c, Ano89f, Ano89g]. Editors [Daa87a, Daa87b, Fur85a, Gar84, Haw82, Hee89, Kes82a, Lee88, Lew85, Lin87, Loc88, Lon85, McI89b, Mit84, Nae86, Som89, Tho86b, Wid86, Wym82, Wym86]. Edn [Car86b, Kar84, Kes83b]. Edward [Ken85]. Eel [RC84, PB86]. eels [Apr87, Mor86]. Effect [BS87, FH83, HW85, MS87, WR85, WH89]. effectiveness [Dic89a, TB88]. Effects [Car88, JK87, LG87, VWN87, Joh88, LPE83, Ste88]. efficacy [BD87]. efficiency [AO87, SH82, YF89, Zal85]. effort [AP86, AO87, GSC89, Han87, PDS86, SBPR89]. egg [ST82]. eggs [AH86, Jør85, PM86]. EH3 [De 86]. electrified [Ste82]. electrofishing [Zal85]. Elmer [Tho89a]. Elsevier [Lee88, Pri84b]. elvers [Apr87]. emergence [HW85]. empirical [HH87]. enclosure [JK87]. encrasicolus [EO89]. end [LSC85]. Energy [Foo82, BK89]. England [Ano86c, Nak85, Pet86, Tho89b, Cru86]. Engraulis [EO89]. enlarged [Tho88]. environment [Sch87, You86]. Environmental [Gjø84b, Wym83, SRRU82]. environmentally [Tan85]. Errata [Ano84c]. Erratum [Ano82c, Ano82d, Ano86e, Ano89h]. errors [LG87]. esculentus [HW85]. Essex [Mit82b, Sae82]. estimate [Lan82, Zal85]. Estimated [GOMH85]. estimates [Dav86, DS89, JK87, LG87]. Estimating [GSC89, Cri83, God89]. Estimation [EO89, Bor89, Wym82]. estimators [Qui85]. estuaries [Mar85]. Estuarine [Mit85, SHS87, Wym86, Fin86, GM82, WW82]. Ethmalosa [Mos88]. ethology [Wym83]. Eugene [All82a]. Europe [Gjø84a, Mit82b, Sae82]. eutrophic [Lee88]. Evaluation [YF89, Cri83, Cus82a, PDS86]. Evan [Wis84]. Everhart [Kes83b]. evidence [GM88]. evolution [Bor86]. expanded [LSM85]. experience [FH83, Ken85]. Experimental [HCG86, GM82, PW86]. experiments [MS85, SF87]. exploitation [Wat82, Som86]. Extended [Fla82].

F [Kes82b, Mit82b, Mit85, Mit88, Paw89, Pil86, Pri86]. F.A.O [Pee84, Tho88]. F.R.G [Paw89, Som89]. Factors [BC83, Mar85]. FAO [Car86a, Dah82b, Fra86, Kar84, Kes83a, Dah82a]. FAO/IDRC [Kes83a]. FAO/Wiley [Fra86]. Farmer [Kir83]. Farming [Wis84]. Farnham [Car86a, Car86b, Cus82a, Dah82b, Hee89, Hop89, Kar84, Kes82d, Kes85b, Lis82, Loc88, Pee84, Pet86, Ruc86, Smi82, Tho88]. Faroe [JO87]. Faroese [HH88]. fauna [GE86]. Fay [Smi83]. Feasibility [DS87]. February [Ano89i]. federal [Kes84b]. Feeding [PC89, NNAS89, TTA82]. Femund [SN89]. fermentation [Fre89]. ferro [Dah82a]. ferro-cement [Dah82a]. Ferto [PB86]. FF [Daa87b]. FF250 [Daa87a]. Fibre [Car86a]. fife [De 86]. Fifth [Joh85]. fight [All83]. figs [All82a, Car86b, Cha82, Kes84a, Kes84d, Kes85b, Lew85, Lon85, Ruc86, Smi83, Som88, Tho88, Thr84, Wal89, Wym84]. Fiji [Joh84]. Fijian [Joh84]. fimbriata [AH86, Mos88]. Finland [JLL84]. Finn [Kes84b]. first [App88, Fag88, Tho89a]. first-prize [Fag88]. Fish [DB89, PJDZ85, AO87, BG85, Bla85, BEKV82, Cro87, Cro89, DS87, DS89, Dic89a, Eve84, Fag88, Foo82, GM82, GW89, Han82, HH87, JK87, JA89, JLL84, Kes82a, Lis82, Mag86, ML88, MAN86, NNAS89, PCS^{+}85, PM86, Pri86, RL89, SHS87, SY84, SD89, Ste82, ST89, UB89a, Van82, Wal89, Wat82, Wym82, Zal85, Ano86c, Bai86, Car86b, Fra86, Haw82, Kes83a, Lin87, Wei85, Wis84]. Fisheries [All89, Bro83, Lan85, Pri84b, Tho86b, Ano88d, Ano89c, BK89, Cha82, Cru86, Cus82a, Fla82, Fur85a, Fur85b, Goe82, Kes82e, Kes82f, Kes85b, Loc88, McI87, McI89a, Mit85, Mit87, Nau85, Pet86, Ruc86, Som87, You86, Kes84d, Mit82b, Mit82a, Paw89, Pil86, Sae82, Tro82, Gjø84a, Nak85, Smi82, Som88, Tho86b]. fishermen [BC83, Joh84]. Fishery [Nau85, Ber88a, Ber88b, Blo87, Bri83, Bri85, Da 85a, Da 85b, DS87, DS89, Edw83, Han87, Kes83b, MS87, MS85, Moh83, Orł89, SN89, SD89, Tan85, Tho89a, Thr85, Gar84, Fin86, Kes84a]. Fishes [Gjø84b, All82a, All89, DM87, GG84, LSC85, Mar85, SJS^{+}89, VBBM89, WW82, Wym83, Gjø84b, Pri84b, Pol83, Thr84, Haw82, McI89b, Smi83, Daa87a, Daa87b]. Fishing [Car86b, Cus82a, Kes82a, Kes82d, Kes83c, Kes85b, Lis82, Loc88, Moh83, Nak85, Pee84, Pet86, Ruc86, Smi82, Suo82, AP86, AO87, BH86, Car83, Car86a, Dah82a, Dah82b, Dav86, De 86, Fer89, Han87, Hop89, Kar84, Ken85, MO86, MS89, PDS86, RC84, Tho88, Suo82]. fit [HH87]. flash [SHS87]. Florida [MM85]. flow [Mat88, SHS87, Ste88]. fluctuations [Orł89]. flume [SF85]. Food [Fra86, Suo82, Wis84, NNAS89, Kes82c, Mit82b]. forms [CCS86]. Forstner [Lew85]. Foster [Car83]. Foundation [Kes85b]. Four [SF87, DG86, Fag88, Mac82]. four-panel [Mac82]. framework [Tro82]. franciscanus [SLH87]. frequencies [SHS87]. Freshwater [Joh85, SY84, Joh85, Pet86]. Fridman [Tho88]. frontier [Cha82]. fry [ST82, TTA82]. Fundamentals [Wid86]. funds [Kes84d]. Further [Van82, SD89]. Fyson [Dah82a, Hop89].

G [Gar84, Gjø84b, Kes85a, Mac83, Mag86, Mit84, Mit87, Pet86, Pil86, Tho88, Wym83, Wym84]. gaboid [MS85]. gadoid [BJD87, DM87]. gadoids [MS87, GS87]. Gadus [FH83, Jør85]. gaffkemia [BD87]. gains [Ull87]. gangs [Tho89b]. Garden [Pet86]. Gary [Wid86]. gayi [CC85a]. gear [AO87, Car83, EW87, GW89, Tho88, Car86a, Kar84]. gear-related [EW87]. Gene [Gjø84b]. general [Bri86]. genetics [KFM87]. genus [All82a]. geographic [Fur85b]. George [Sae82, BC83]. Gerhard [Car86a, Kar84]. Gesellschaft [Som89]. Ghana [Bla85, Van82]. gigas [App88]. gill [Bor89, DS89, Jen82, Mat88, SN89, SF85]. gill-net [Bor89, DS89, SN89]. gillnet [BC88, DS87, Dic89a, Dic89b, SD89]. gillnets [HCG86, Mar85, Ste88, TB88]. given [Bai86]. giving [All82b]. gladius [TT89]. Glasgow [Pol83]. Glossary [Car83]. Gnaiger [Lew85]. Goldfish [Wis84]. Goldman [Joh85]. gooseneck [Ber88a]. Goulding [Cha82]. Grand [BC89]. Grant [Eve84, Kes84d, Mit82a, Mit84, Som87, You86]. grass [HOE89]. Gratzek [Wis84]. Great [Cus82a, Dah82a, Dah82b, Kes82a, Kes82d, Lis82, Mit82b, Sae82, Smi82, Suo82, Goe82]. Greece [Pap84, PC89]. Greenland [BC89]. grenadier [Atk89]. grey [TTA82]. Groman [Lan85]. Grosvenor [Tho86b]. ground [Ane89]. Growth [GM87, GM88, HH88, Mos88, App88, BCE86, Cro89, DG86, DGMO88, EO89, Hal84, HH87, HW85, LG87, MD87, MM85, MWS83, NR87, Pap84, PB86, SGB87, TT89, All89]. Gt [Car86a, De 86, Fra86, Hee89, Kar84, Kes85b, Ruc86]. GTZ [Som89]. GU12 [Hee89]. Guichenot [CC85a]. Gulf [Dav86, Pap84, You86, AH86, ANM87, GOMH85]. Gulland [Fra86]. Gunn. [Atk89]. Guraya [Wal89]. gurnard [Pap84].

H [Cus82a, Eve84, Lew85]. haddock [EW87]. Hague [All82a, Cha82, Gjø84b, Wym83]. Haines [Nau85]. hake [BA82, CC85a, LSM85]. Halapua [Joh84]. Halibut [BC89, HT86, HH88]. Haliotis [MS89]. Hall [Pol83]. Hamabe [Pee84]. Hammar [NNAS89]. Hamuro [Pee84]. Handbook [Wis84, Da 85a, Da 85b]. handline [Ber88b]. handling [WH89]. Hanover [Tho89b]. Hants [Hee89]. Hara [Pri84b]. hard [VWN87]. hard-bottom [VWN87]. hardback [Bai86, Dah82a, Joh85, Mit82b]. hardcover [Ano86c, Bor86, Som89]. Hardy [Lis82]. harengus [Ane89, KFM87]. Hariati [All89]. Harry [Kes83b]. Hart [Gjø84a]. harvest [SRRU82]. harvested [SLH87]. hasselti [SY84]. Haswell [HW85]. hatcheries [FS89]. hatchery [ST89]. hauls [War89]. Hawaii [Thr84]. Hayes [Smi82]. Headline [Mat88, Ste88]. Hearing [Smi83]. Heemstra [McI89b]. Heidelberg [Fin86, Kes84b, Kir86, Lew85, Lon85, Mit85, Wym86]. Heidelberger [Nae86]. height [Mat88]. heights [Ste88]. held [Wym83]. Helfman [Gjø84b]. Helm [Ano86c, Gjø84a, Lin87, Mit88, Pri89]. Hemisphere [Wid86]. Herrin [De 86]. Herring [Ane89, De 86, HCG86, KFM87, RP82, RLH88]. herrings [SJ86]. high [GW89, PW85, PW86]. hippoglossoides [BC89]. Hippoglossus [HT86, HH88]. Histology [Lan85]. Historical [BWL87]. history [BJD87, De 86, DM87]. Holmgren [Lin87]. Homarus [BD87, DB89]. Home [WB82]. hook [FH83]. hooks [Ber88b]. Hopkins [Mit87]. Horwood [Pri89]. Hungary [PB86]. Hureau [Daa87a, Daa87b]. hydroacoustic [JLL84, Orł89]. hydroacoustical [Wym82]. Hydrobiology [Cha82]. hypersaline [RC84].

I.C.L.A.R.M [Bai86]. Ian [Ruc86, Tro82]. iced [SJS^{+}89]. Iceland [JO87]. ichthyoplankton [LT89]. ICLARM [All89, Cus82b, Fur85a, Fur85b, Gar84, Haw82, Paw89, Som89, Thr85, Tro82].

Identification [CCS86, All82b]. IDRC [Kes83a, Kes83a]. IDRC-198e [Kes83a]. II [Daa87b, SD89]. ill [Lin87, Loc88, Mit88, Som87]. illus [All83, Car83, Cus82a, Dah82a, Dah82b, Kes82a, Kes82d, Kes83b, Lis82, Pee84, Smi82, Suo82, Wei85]. illustrated [Bai86, De 86, Eve84, Nau85]. illustrations [Car86a, Kar84, Wis84]. IMARPE [Som89]. immature [GOMH85]. Impact [AO87, SN89]. Impacts [Nau85, Blo87]. implications [Wat82]. important [All89, SJS^{+}89]. importation [MO86]. improved [Ull87]. include [Fre89, Tho89a]. including [Tho89a]. index [Ano82a, Ano82f, Ano85a, Ano85f, Ano86b, Ano86i, Ano87a, Ano87g, Ano88h, Ano89b]. India [DS88, Kes84a]. Indian [Ber88b]. indices [SGB87]. Indo [Thr84]. Indonesia [All89]. Indonesian [All89, Paw89]. induced [Tan85]. industry [Fer89]. infection [CC85a]. inferred [MYH86]. influence [SSR89]. influenced [SHS87]. influencing [Mar85]. inhabiting [CCS86]. inkwell [SH82]. inshore [ML88, Mos88]. Institute [All89, Joh84, Kes84c, Kir83]. Instituto [Som89]. instrument [Goe82]. intensities [GW89]. interaction [PC89]. interactions [MS87]. Interannual [You86, FS89]. International [All89, Bai86, Bor86, Haw82, Joh85, Ken85, Kes82b, Kes83a, Kir83, Nau85, Paw89, Som89, Kes82e, Kes84c, Tho89a, Mit82a]. Interscience [Fre89]. introduction [DS87, Moh83]. Investigaciones [Kes84c]. investigates [Dah82a]. investigating [SF87]. investigation [Ste82]. investment [BK89]. Iraq [NNAS89]. Ireland [Bri83, Bri85]. Irish [Cro87, Cro89]. irrigation [ST82]. irroratus [BCE86]. ISBN [All82a, All83, All89, Ano86c, Bai86, Bor86, Car83, Car86b, Cha82, Cus82a, Da 85a, Da 85b, Daa87a, Daa87b, Dah82a, Dah82b, De 86, Fin86, Fra86, Fre89, Gar84, Gjø84a, Gjø84b, Haw82, Hee89, Hop89, Joh85, Kar84, Kes82a, Kes82c, Kes82d, Kes83a, Kes83b, Kes83c, Kes84b, Kes85b, Kir86, Lan85, Lee88, Lew85, Lin87, Lis82, Loc88, Lon85, Mac83, Mag86, McI89b, Mit82b, Mit84, Mit85, Mit87, Mit88, Moh83, Nae86, Nak85, Paw89, Pee84, Pet86, Pol83, Pri84b, Pri86, Pri89, Ruc86, Sae82, Smi82, Smi83, Som86, Som87, Som88, Som89, Suo82, Tho86b, Tho88, Tho89b, Tho89a, Thr84, Thr85, UB89b, Wal89, Wei85, Wid86, Wis84, Wym83, Wym84, Wym86, Car86a, Bro83]. ISBN-0-85238-123-9 [Car86a]. Island [PC89, Sae82]. Islands [JO87, MD87, Ken85]. Isles [KFM87, Kes85b]. ISSN [All89, Cus82b, Fur85a, Fur85b, Ken85, Kes85a, Lan85, McI89a, Tro82]. issue [Kes82e]. issues [McI89a]. Ithaca [Kes83b, Nau85]. iv [Kes84d]. ix [All82a, Kes84b, Pol83, Tro82].

J [Car83, Daa87a, Daa87b, Fin86, Fra86, Gjø84a, Gjø84b, Haw82, Hop89, Kes82a, Kes82b, Lis82, Lon85, Mit84, Mit85, Nae86, Pri84b, Pri86, Pri89, Tho88, Thr84, Thr85, Wym82, Wym83, Wym86]. J.-C [Daa87a, Daa87b]. jackass [WHW88]. Jackson [Fla82]. Jakarta [All89]. James [Kes82c, Kes85b, Mit88, Tho89b]. January [Ano88e, Gar84]. Japanese [Ken85]. japonicus [TL83]. Jersey [Wym86]. Jigging [Pee84]. Johannes [Kes83c]. John [Dah82a, De 86, Fra86, Mit87, Wis84]. Johnson [Bor86, Nau85, Tho86b]. joint [Ken85, Mit82a]. Jon [Kes85a]. Jorge [Eve84]. Jr [Mit84]. July [Ano86g, Ano89j]. June [Ano83c, Ano89l, Wym82]. Junk [All82a, Cha82, Gjø84b, Wym83]. Juridicas [Kes84c]. jurisdiction [Fla82]. juvenile [BCE86, MWS83, SGB87].

Kaczynski [Mit82a]. Karger [Wal89]. Kariba [ML88]. Kastellorizo [PC89]. Kattegat [RP82]. Keen [Tho89a]. Kennish [Wym86]. Kent [Ano86c, Gjø84a, Pri89, Sae82]. Kerala [DS88]. Kingdom [Car83]. kisutch [Joh88]. Klust [Car86a, Kar84]. Kokanee [ST82]. Kondrashova [Mag86]. Konferenzen [Som86]. Kuwait [Kir83, AH86, Kir83]. Kuwaiti [AH84].

L [All89, Daa87a, Daa87b, Fin86, Hee89, Smi82, Tho86b, Tho88, Thr85, Wym83]. L. [Ane89, App88, Apr87, Bor89, Bri86, CCS86, Cro87, Cro89, EO89, FH83, Jør85, LPE83, Mor86, NC89, Pap84, PB86, SN89, TT89]. Laboratories [Wym82]. laboratory [HW85]. Laevastu [Cus82a, Smi82]. Lagoon [Kes83c, RC84]. Lake [Van82, AP86, Bor89, RL89, WB82, JLL84, ML88, PB86, SN89, TB88]. lake-specific [RL89]. lakes [UB89a, Lee88, UB89b]. Lammas [De 86]. Lane [Tho86b]. Langen [Paw89, Som89]. Lanka [AP86, DG86, DS87, DS89, SD89]. large [FS89, Hee89, TBH85, UB89b]. large-capacity [FS89]. Larkins [Cus82a]. Larry [Tho86b]. Larvae [Thr84, Jør85, PM86]. Larval [SGB87]. Lates [Dav86]. lavaretus [SN89, SVQ85]. law [Bor86]. LCCCN [Nau85]. Leach [MS89]. learning [ST89]. Least [WW82]. Least-cost [WW82]. Lecture [Fin86, Mit85, Wym86]. Lee [Mit87]. legal [Mit85]. legislacion [Kes84c]. Leis [Thr84]. length [Atk89, NNAS89]. Lenz [Lon85]. Leslie [Cri83, San88]. lethal [BG85]. Level [Pol83, AO87, MS85]. levels [All82b]. LeVieil [Mit82a]. Lewis [UB89b]. Lic [Kes82b]. Life [BJD87, ST89, DM87, Lan82, TL83]. light [GW89, PCS^{+}85, SHS87]. light/air [PCS^{+}85]. light/air-bubble [PCS^{+}85]. light/bubble [SHS87]. light/bubble-curtain [SHS87]. Lindquist [Gjø84b]. Line [Qui85, Dah82b]. linear [HH87]. lion [BA82]. LiPuma [Ken85]. liver [BG85]. Living [All89, Bai86, Haw82, Ken85, Paw89, Som89]. lobster [BD87, BC83, DB89, NC89, PDS86, Tho89b]. lobsters [MR85, SH82]. local [Dic89a]. locomotion [Wei85]. logging [BEKV82]. logistic [Jen82]. London [All83, Kes82c, Kes83b, Kes83c, Mit88, Pol83, Tho89b, Wei85, Wym84]. Long [Fag88, Pet86]. Long-term [Fag88]. Lophius [Cro87, Cro89]. lore [Kes83c]. low [GW89, PW85, PW86]. low-aspect-ratio [PW85, PW86]. Lower [Hee89]. Ltd [Dah82a, Dah82b, Hop89, Kar84, Kes82a, Lis82, Loc88, Smi82, Suo82, Tho88]. lucerna [Pap84]. Lury [San88]. Lutz [Wym86].

M [All89, Da 85a, Da 85b, Daa87a, Daa87b, Dah82b, Haw82, Hee89, Kes82d, Kir86, McI89b, Mit84, Som86, Tho89b, Thr84, Wid86, Wym86]. M.-L [Daa87a, Daa87b]. M.C.C [Bai86]. MA [Wym82]. maccoyii [Tho86a]. mackerel [LPE83]. macropterus [WHW88]. made [AP86]. Magnuson [Haw82]. Maine [Tho89b]. Makati [Bai86, Paw89, Som89]. Malawi [TB88]. Malaysia [SY84, Wat82]. Malcolm [Moh83]. Malha [Ber88b]. man [AP86, Cha82]. man-made [AP86]. Management [All89, Bai86, Bro83, Haw82, Ken85, Kes84a, Paw89, Som89, Cus82b, Gar84, GS87, Hee89, Kir86, RC84, TBH85, Ull87, Wat82, Mit87, Cus82a, Fin86, Pet86, Pri89]. Managing [Kes84b]. Mangel [Kir86]. Manila [Bai86, Cus82b, Fur85a, Fur85b, Gar84, Haw82, Ken85, McI89a, Paw89, Thr85, Tro82]. Manilla [Som89]. Manual [Kar84, Pee84, Tho88, Bai86, Fra86]. Manuals [Car86a, Suo82]. Mar [Som89]. Marahudin [Paw89]. March [Ano86f]. Margaret [McI89b]. Margetts [Car83]. Marine [All89, Joh84, Mit82a, Hee89, JA89, Kes83c, Kes84b, Paw89, Som86, Tho89a, Wym82, Cus82a, Kes85a]. Maritime [Bro83]. market [Han82]. marketing [Mit88, Ruc86]. Marking [Kir83]. Marsh [NNAS89]. Marshall [Pol83]. Martyshev [Pil86]. maruadsi [ZW88]. Mason [Kes85b]. Massachusetts [Nae86]. materials [Kar84, MO86, MR85]. Mathematical [Pak89]. Maturation [GM87, HH88]. maturity [HT86]. maximum [All82b, Fag88, GSC89]. maximus [Pri84a]. May [Ano87d, Ano88f, Ano89k, Haw82, Som86]. MC [Paw89, Som89]. McDonald [Tho89a]. McGraw [Kes84d]. McHugh [Fin86]. McKeown [Ano86c]. MD [Lan85, Mit87, Nau85, Tho86b]. meal [Han82]. Mean [San88]. measurement [Fer89, MS89]. Measurements [SF85, Ste88]. measures [Blo87]. mechanised [DS88]. media [BG85]. Mediterranean [Daa87a, GG84, TL83, WR85, Daa87b]. Meeting [Wym82]. Melicertus [TL83]. Meltzoff [Ken85]. Memoria [Kes84c]. menhaden [Blo87]. Merah [SY84]. Merluccius [CC85a]. Merriman [Nae86]. mesh [Bri86, Car88, LT89, LSC85]. Mesta [PJDZ85]. metabolism [BG85]. Metanephrops [WD87]. method [CC85b, Cri83]. methods [Car86b, JLL84, RLH88, San88, Wym82, Fra86]. Metro [Cus82b, Fur85a, Fur85b, Paw89, Som89, Thr85, Tro82]. Mexico [Kes82b, Kes84c, GOMH85]. Michael [Cha82]. Michio [Pee84]. microbiology [Fre89]. Micronesia [Kes83c]. Migration [Dav86, TTA82, JO87, Mor86, Ano86c, Pri86]. Miguel [Fur85a, Fur85b]. Mines [Fur85a]. miniature [GM82]. minimise [MS85]. minor [DS87, DS89, SD89]. Mitson [Nak85]. mm [Cus82a, Dah82a, Dah82b, Lis82, Smi82, Suo82]. mobility [Fur85b]. model [Dic89b, DGMO88, Hal84, Han82, Pak89, SRRU82, Tan85]. models [Han87, HH87]. Modification [Tan85, CC85b]. Molecular [Eve84, Fre89, Wal89]. monitored [JLL84]. monofilament [SSR89]. Monograph [Lan85]. Monographs [Wal89]. Moreton [WH89]. morhua [FH83, Jør85]. Morone [SRRU82]. mortality [All89, App88, CC85b, Dav86, GM88, HW85, LG87, LPE83, MS89, Mos88, WHW88]. Morten [Lee88]. morwong [WHW88]. Mototsuqu [Pee84]. moulting [HW85]. Mountain [PJDZ85]. movement [Dic89a]. Movements [MYH86, HOE89]. MRMS [Bro83]. Muir [Mit88]. mullet [TTA82]. multifilament [SSR89]. multiple [Hal84]. Multispecies [Gar84, Cus82b]. Munro [Thr85]. Murphy [Gar84]. Murray [Smi82]. muscle [JA89]. mutton [MM85].

N [Fur85a, Mag86, Pol83, Smi83]. N.S.W [Thr84]. national [Fla82, Ken85]. native [VBBM89]. natural [CC85b, GM88]. negligible [CC85b]. Nemadactylus [WHW88]. Nephrops [Bri83, Bri85, Bri86, Car88, MS85, NC89]. nerka [ST82]. Net [Bri83, Bor89, DS89, Jen82, MS85, SN89]. Netboy [All83]. Netherlands [All82a, Da 85a, Da 85b, Gjø84b, Som88]. nets [LT89, Mat88, SF85]. netting [Fer89, SSR89, Kar84]. News [Car83, Car86a, Car86b, Cus82a, Dah82a, Dah82b, Hop89, Kar84, Kes82a, Kes82d, Kes85b, Lis82, Loc88, Moh83, Nak85, Pee84, Pet86, Ruc86, Smi82, Suo82, Tho88]. Nicolson [Kes82c]. Nielsen [Daa87a, Daa87b, Tho86b]. Nigeria [MO86]. Nigerian [Mos88]. Nilsson [Lin87]. Niño [Joh88]. nitrogen [Jør85]. No [Bro83, Fur85a, Fur85b, Haw82, Kes82b, Lan85, McI89a, Pri84b, Thr85, Tro82]. Noakes [Gjø84b, Wym83]. nocturnal [HW85]. non [HH87]. non-linear [HH87]. Norbert [UB89b]. Normal [Wym83]. North [Daa87a, Daa87b, Nau85, Cro89, Atk89, Cro87, GS87, HT86, LSC85, MS87, RP82, Lon85]. North-eastern [Daa87a, Daa87b]. North-Western [LSC85]. Northeast [DM87, GM87, NR87, SL87, Ull87]. Northeastern [Nau85]. northern [BC89, Bri83, Bri85, Jak87]. Northwest [BC89, GOMH85, WD87]. norvegicus [Bri86, NC89]. Norway [Jak87, NC89, SN89]. Norwegian [GM87, HT86, JO87]. note [Ano85d, Ano85e]. Notes [Mit85, Wym86, Fin86]. Nova [Bro83, Bro83]. November [Ano87e, Ano89m, Kir83]. Novikov [Da 85a, Da 85b]. Number [Kes85a, JLL84, San88]. NY [Nau85]. nylon [FS87].

O [Paw89, Som89]. O. [Joh88]. Oberservations [Nae86]. Objective [Fer89]. observation [Van82]. Observations [All82b, Bla85, LS89]. observed [GG84]. Occupational [Fur85b]. Occurrence [Cro87]. Ocean [Ber88b, Mit84, Kes84b, Mit82b]. oceanography [Smi82]. oceans [GS87]. October [Kir83, Wym83]. off [App88, BYNB87, DP89, MS89, MWS83]. Ogura [Pee84]. Oncorhynchus [Fag88, Joh88, ST82]. one [McI89a]. oogenesis [Wal89]. Optimal [Han87, God89]. optimization [BH86]. Orange [TBH85]. orders [Tho89a]. Oregon [Kir86, Joh88, MWS83]. Organization [Suo82]. orientation [NC89]. Osteochilus [SY84]. other [Tho89a, WD87]. Otolithes [AH84]. otoliths [SVQ85]. Ottawa [Kes83a]. otter [GM82, PW85, PW86, SSR89]. Ownership [Tho89a]. oxygen [Lew85]. oxytetracycline [BD87]. oyster [Hal84].

P [Car83, Daa87a, Daa87b, Fre89, Gjø84a, Kes84a, Kes84b, Paw89, Som89, Tho88, Wym84, Kar84, Moh83]. P.O [Bai86]. Paasivesi [JLL84]. Pacific [Joh84, Sae82, Thr84, BA82, DM87, GS87, HCG86, Kes84d]. package [Goe82]. packing [Car83, Kar84, Lis82, Pee84, Smi82]. pages [Lee88, Ano82e, Ano83c, Ano83d, Ano84d, Ano84e, Ano85c, Ano86h, Ano86g, Ano86f, Ano87f, Ano87d, Ano87e, Ano88e, Ano88f, Ano88g, Ano89i, Ano89l, Ano89j, Ano89k, Ano89m]. Pagrus [PC89]. pair [Edw83]. Palau [Kes83c]. palliative [Fla82]. Palomares [All89]. Pampus [AH84]. panacea [Fla82]. Pandalus [BS87]. panel [Mac82]. Panulirus [BC83]. paper [Thr85]. paperback [Cus82b, Fla82, Gar84, Haw82, Kes84b, Mit82b, Pol83, Tro82, Thr85]. paperback/US [Mit82b]. Papers [Joh85]. Parakrama [AP86]. parameters [EO89]. parents [Jør85]. Paris [Daa87b]. Parrish [Loc88]. part [WR85]. particular [Tan85]. pathways [Fre89]. patterns [Dav86, WB82]. Pauly [All89, Bai86, Cus82b, Fur85a, Gar84, Som89]. Pecten [GM88, SGB87]. pelagic [JLL84, SN89, WW82]. penaeid [All89]. Penaeus [HW85, TL83]. perch [SRRU82]. performance [DS88, EW87]. period [GM88]. periodicity [WD87]. Personal [Cru86]. Perspectives [All82a]. Peru [Som89]. Peruvian [Som89]. Pesca [Kes84c]. pesquero [Kes82b, Kes82b, Kes84c]. Peter [De 86]. Petrocelli [Wid86]. Philippines [Bai86, Cus82b, Fur85a, Fur85b, Gar84, Haw82, McI89a, Paw89, Som89, Thr85, Tro82]. Phillip [GM88, McI89b, SGB87]. photography [GG84]. Phycis [PC89]. Physical [LPE83, Sch87]. physiological [BYNB87, Lew85]. physiology [Lin87]. Pilleri [Wym84]. Pioneer [Kes84a]. piscatorius [Cro87, Cro89]. Pitcher [Gjø84a]. planning [BK89]. plates [All82a, Kes84a, Kes85b]. platyceros [BS87]. Platz [Nae86]. plumb [GE86]. plus [Joh84, Kar84, Pee84]. polar [Bor86]. Polarographic [Lew85]. pole [Dah82b]. policies [MO86]. Polish [Orł89]. politics [Mit82b, Sae82]. Pollicipes [Ber88a]. pollock [ANM87, BWL87, LG87]. polymerus [Ber88a]. polymorphic [SN89]. Pomadasys [AH84]. Pond [Pil86]. Popper [Smi83]. Population [Pri89, WHW88, ZW88, All82b, Bai86, BC88, Cri83, GR82, KFM87, LSM85, Pak89, San88, SN89, WR85]. populations [Bla85, Qui85, Wym82]. Port [GM88, SGB87]. Portland [Mit88]. Portuguese [Car88]. post [Car83, Kar84, Lis82, Pee84, Smi82]. post/packing [Car83, Kar84, Lis82, Pee84, Smi82]. postage [Joh84]. Postal [Kes82b]. Postfach [Nae86]. pot [SH82]. potamodromous [MYH86]. Potential [Ber88a, Ull87, EW87, Mos88, SGB87]. pots [YF89]. pp [All82a, All83, All89, Ano86c, Bai86, Bor86, Bro83, Car83, Car86a, Car86b, Cha82, Cus82a, Cus82b, Da 85a, Da 85b, Daa87a, Daa87b, Dah82a, Dah82b, De 86, Eve84, Fin86, Fla82, Fra86, Fre89, Fur85a, Fur85b, Gar84, Gjø84a, Gjø84b, Haw82, Hee89, Hop89, Joh85, Kar84, Ken85, Kes82a, Kes82c, Kes82d, Kes83b, Kes83c, Kes84a, Kes84b, Kes84c, Kes84d, Kes85a, Kes85b, Kir86, Lan85, Lew85, Lin87, Lis82, Loc88, Lon85, Mac83, Mag86, McI89a, McI89b, Mit82b, Mit82a, Mit84, Mit85, Mit87, Mit88, Moh83, Nae86, Nak85, Nau85, Paw89, Pee84, Pet86, Pil86, Pol83, Pri84b, Pri86, Pri89, Ruc86, Sae82, Smi82, Smi83, Som86, Som87, Som88, Som89, Suo82, Tho86b, Tho88, Tho89b, Tho89a, Thr84, Thr85, Tro82, UB89b, Wal89, Wei85]. pp [Wid86, Wis84, Wym82, Wym83, Wym84, Wym86, You86, Wym82]. prawn [BS87, HW85, SSR89, WH89]. prawns [YF89]. Precipitation [Nau85]. predation [BA82]. Predators [Gjø84b]. predict [SGB87]. Prediction [RL89]. Preliminary [BYNB87, ML88]. prepared [Tho89a]. Press [All83, Daa87a, Daa87b, Fla82, Hee89, Kes83b, Kes83c, Mit82b, Mit84, Mit87, Mit88, Sae82, Tho89b, Thr84, Wei85, Wym84]. Prey [Gjø84b]. Price [Haw82, Bai86, Bor86, Car86a, De 86, Fin86, Fur85a, Fur85b, Gar84, Gjø84a, Gjø84b, Ken85, Kes84b, Kes84d, Kir86, Lan85, Nae86, Nak85, Pee84, Pil86, Pol83, Pri84b, Ruc86, Som89, Tho86b, Thr85, Wym84, You86]. Principles [Kes83b]. Printed [Wym83]. prize [Fag88]. Proceedings [Gar84, Haw82, Nau85, Wym83, Kir86, Kir83]. processing [BEKV82, Mag86]. production [FS89, GSC89, PJDZ85]. productivity [Tho89a]. products [Moh83]. program [Kes84b]. programmable [Bai86]. proposed [Blo87]. Pseudotolithus [DP89]. Publ [Wym82]. Publication [Eve84, Fre89, Mit84]. Published [Dah82b, Gjø84a, Joh84, Mac83, McI89a, Nau85, Gar84]. Publisher [Ano85e, Ano85d]. Publishers [Da 85a, Da 85b, De 86, Kes84a, Lin87, Pri84b, Som88, UB89b]. Publishing [Bor86, Gjø84a, Joh85, Lee88, Mit82b, Sae82, Tho89a, Wid86, Wis84]. Puerto [App88, MD87]. Pulvenis [Mit85]. Purves [Wym84]. Pushnet [TNW89].

Q [Pol83]. quality [DB89]. quantitative [Cus82a]. queen [Kes85b, MD87, App88]. Quensel [LS89].

R [Car83, Haw82, Joh85, Kes83c, Lis82, Loc88, Mit85, Nak85, Smi83, Som86, Tro82, Wei85, Wid86, Wym86]. radiotelemetric [HOE89]. Raiffenstrasse [Paw89, Som89]. rain [Nau85]. rain/fisheries [Nau85]. Rand [Wid86]. range [WB82]. Rao [Kes84a]. Rapporteur [Sch87]. rate [GM88, SHS87]. rates [GOMH85, JO87, WHW88, WH89]. ratio [PW85, PW86]. Raymond [Nau85]. reactions [GW89, Ste82]. reappraisal [ML88]. recapture [BC83]. recruit [BH86, LG87]. Recruitment [Kir83, SLH87, All82b, All89, SGB87, Tan85]. red [DGMO88, GOMH85, SLH87]. Reef [Thr84, Goe82, Thr85]. reference [Tan85]. references [Car86a]. refrigeration [Mag86]. refs [Car86b]. regard [Bor86]. regime [Mit85]. Reginald [Pri86]. region [Mit85]. regression [Hal84]. regulation [GR82]. Regulations [MO86]. Reinhardtius [BC89]. Reinhart [Haw82]. related [Dic89a, EW87]. relation [LS89, NR87, NC89, SRRU82]. relations [RP82]. relationships [Atk89, NNAS89, Zal85]. relative [BC89]. release [DB89]. Releasing [Kir83]. remote [Pri84a]. Rennis [Thr84]. Replica [Sae82]. Report [Fur85a, Fur85b, Mit82a, BYNB87, Sch87]. Reports [All89]. reproduction [App88, BH86, SL87]. reproductive [AH84, Tho86a]. Research [All89, Kes82a, Kes83a, Kir83, Ano88d, Ano89c, Goe82, Kes82e, Kes82f, Loc88, McI87, Tro82, VWN87]. Reservoir [SY84, Bla85, TBH85, VBBM89]. reservoirs [DS87, DS89, SD89]. Resource [Bro83, BK89, Kes85a, Kir86]. Resources [All89, Bai86, Haw82, Joh84, Ken85, Mit82a, Paw89, Som89, BJD87, Bor86, DS87, DS89, Kes84b, SD89, Mit84, Tho89a, Thr85]. Responses [PCS^{+}85, SHS87]. results [Orł89, DS89]. returned [BC83]. Review [All82a, All83, All89, Ano86c, Bai86, Bor86, Bro83, Car83, Car86a, Car86b, Cha82, Cus82a, Cus82b, Da 85a, Da 85b, Daa87a, Daa87b, Dah82a, Dah82b, De 86, Eve84, Fin86, Fla82, Fra86, Fre89, Fur85a, Fur85b, Gar84, Gjø84a, Gjø84b, Haw82, Hee89, Hop89, Joh84, Joh85, Kar84, Ken85, Kes82a, Kes82b, Kes82c, Kes82d, Kes83a, Kes83b, Kes83c, Kes84a, Kes84b, Kes84c, Kes84d, Kes85a, Kes85b, Kir83, Kir86, Lan85, Lee88, Lew85, Lin87, Lis82, Loc88, Lon85, Mac83, Mag86, McI89a, McI89b, Mit82b, Mit82a, Mit84, Mit85, Mit87, Mit88, Moh83, Nae86, Nak85, Nau85, Paw89, Pee84, Pet86, Pil86, Pol83, Pri84b, Pri86, Pri89, Ruc86, Sae82, Smi82, Smi83, Som86, Som87, Som88, Som89, Suo82, Tho86b, Tho88, Tho89b, Tho89a, Thr84]. Review [Thr85, Tro82, UB89b, Wal89, Wei85, Wid86, Wis84, Wym82, Wym83, Wym84, Wym86, You86, Bri86, DM87, MAN86]. Reviews [Som89, Thr85, Tro82]. revised [Tho88]. revolution [SJ86]. rhythms [TTA82]. Ricker [Tan85]. Rico [App88, MD87]. Riemann [Lee88]. Rila [PJDZ85]. River [PJDZ85, TBH85, All83, TTA82, Wat82]. Riverine [Mor86]. rivers [TNW89, Zal85]. roach [Bor89]. Road [Hee89]. Robin [Pet86]. rock [BCE86, BC83]. role [Kes84b]. ropes [Car86a]. Rotterdam [Da 85a, Da 85b, Mag86, Pil86, Som88]. roundnose [Atk89]. Roy [Fla82]. Rs [Kes84a]. rubra [MS89]. Rudolphi [Cro87]. Rule [ST82]. rupestris [Atk89]. Russak [Kes85a]. Russell [Mac83]. Russell-Cargill [Mac83]. Rutilus [Bor89].

S [All89, Car83, Gjø84b, Joh84, Ken85, Kes84a, Kir83, Loc88, Thr84, Wal89, You86, ZW88]. s.l [SN89]. Safety [BD87]. Sainsbury [Fre89]. saithe [JO87]. salar [CCS86]. Salmiya [Kir83]. Salmo [CCS86, MYH86]. Salmon [FS89, Joh88, All83, CCS86, Fag88, GR82, ST82, Mit88]. Salmonid [All82a]. salmonids [MWS83]. saltmarsh [LS89]. Salvelinus [All82a]. Sam [Wid86]. same [CCS86]. Samokov [PJDZ85]. sample [GM82]. sampler [PM86]. sampling [EW87, GE86, JK87, LT89, TNW89, UB89a]. Samudra [AP86]. San [Fur85a, Fur85b]. sanctuaries [Kes84b]. Sarah [Ken85]. Sarawak [Wat82]. sardine [AH86]. Sardinella [AH86, DG86]. Sardinia [RC84]. satellite [Pri84a]. Saya [Ber88b]. scale [BK89, DS89, MYH86, Fur85a, Fur85b]. scales [SVQ85]. scallop [GM88, SGB87, Kes85b]. scan [Mac83]. Schink [Kes84d]. Schmidtke [UB89b]. schooling [Qui85]. sciaenid [DP89]. Science [All82a, Lee88, Pri84b, Kes82a, McI89a, Kes83b]. sciences [Som87]. Scientific [Kes82f, Kir83, Paw89, Som89]. Scomber [LPE83]. scombrus [LPE83]. scope [McI87]. Scotia [Bro83]. Scotland [Loc88]. Scottish [SH82]. Sea [Eve84, Kes84d, Mit82a, Mit84, You86, BA82, Bor86, BC83, Kes83a, Lon85, McI89b, Mit82b, SLH87, Ane89, BWL87, BJD87, Cro87, Cro89, EW87, EO89, JA89, KFM87, MS87, NR87, RP82, TT89, WR85, ZW88, You86, Kes82c, Kes82d]. seals [Lan82]. season [HH88]. Seasonal [Orł89, TTA82, UB89a, HH87, VBBM89]. Seattle [All83, Eve84, Mit84, You86]. Second [Fre89]. sei [Pri89]. seine [TNW89]. selection [Bri86, Jen82, LT89]. Selectivity [Ber88b, Bor89, BC88, Bri83, LSC85]. self [PM86]. self-adjusting [PM86]. senegalensis [DP89]. sensing [Pri84a]. sensors [Lew85]. September [Ano83d, Ano84e, Ano88g]. Series [Fra86, Pol83, GSC89]. Service [Bro83, Hee89]. set [Mat88, Ste88, UB89b]. settlement [BCE86, SGB87]. sex [EO89]. sexual [HT86, HH88]. shallow [GM82]. shark [Pri84a]. sharks [GG84]. sharpeyi [NNAS89]. Shaw [Mit88]. Sherman [Hee89]. ship [BEKV82]. ship-borne [BEKV82]. Shipboard [Mag86]. shipping [Tho89a]. short [War89]. should [Tho89a]. Shrimp [Kir83, WR85]. shrimps [All89]. side [Mac83]. Sierra [Kes82b]. simulated [GR82]. simulation [Dic89b, Som87]. simulations [Cru86]. Simultaneous [BC88, Pri84a]. single [RLH88]. single-beam [RLH88]. Singleton [Fre89]. situ [RLH88]. size [Car88, Cri83, EO89, Fag88, HT86, Mar85, Pak89, Zal85]. size- [HT86]. size-structured [Pak89]. Skagerrak [RP82]. skills [ST89]. sled [PM86]. Small [Pee84, BK89, Bla85, Bor89, Hop89, UB89a, YF89, Fur85a, Fur85b]. small-scale [BK89, Fur85a, Fur85b]. Smith [De 86, McI89b, Pri86, Tro82, McI89b]. snapper [MD87, MM85]. soak [BS87]. sobre [Kes84c]. Social [ST89]. Society [Lan85, McI89a, Nau85, Tho86b]. soft [Fin86, Mit85, Wym86]. Softcover [Pet86, Bor86]. sole [LS89]. Solea [LS89]. Solomon [Ken85]. Some [BK89, Mar85, PW85, LSC85, MS87, Pap84, WW82]. Son [Pol83]. sonar [Mac83, RLH88, Suo82, Nak85]. Sondergaard [Lee88]. sound [Smi83]. sounding [Suo82]. Sources [JK87, EW87]. South [Joh84, Mac83, TBH85, TL83, WR85, ZW88]. south-eastern [TL83, WR85]. southeastern [MS89, Fag88, SBPR89]. southern [RC84, Tho86a, SLH87]. Sowerby [Ber88a]. spatfall [SGB87]. Spatial [VBBM89]. spawning [Ane89, CCS86, God89, HH88, Pak89]. spearfishermen [SBPR89]. special [Bor86, Wym83]. species [DG86, DP89, JA89, MS87, PC89]. specific [Dav86, RL89]. specimen [Zal85]. speeds [WW82]. spiny [MR85]. spoilage [SJS^{+}89]. sponges [VWN87]. sport [Fag88]. sport-caught [Fag88]. spp [AH84]. sprat [RLH88]. Spring [ZW88]. Springer [Fin86, Kes84b, Kir86, Lew85, Lon85, McI89b, Mit85, Nae86, Pri86, Smi83, Som86, Wym86]. Springer-Verlag [Fin86, Kes84b, Kir86, Lew85, Lon85, McI89b, Mit85, Nae86, Pri86, Smi83, Som86]. Squid [Pee84]. Sri [AP86, DG86, DS87, DS89, SD89]. St [De 86]. staff [GE86, Kes82a]. Stark [Wym82]. State [Kir83, DS88]. States [Kes84b]. Static [BH86]. Station [Kes82a]. statistical [SRRU82]. status [BYNB87]. steelhead [All83]. Stefan [Lin87]. Stephen [De 86]. stochastic [Han87]. Stock [SJ86, DP89, Fra86, Fur85a, GR82, Tan85, Ull87, ZW88]. stocking [Apr87].

Stocks [Gar84, Car88, ML88, RP82, Van82, Cus82b]. storage [SJS^{+}89]. Strait [WHW88, BC89, BYNB87]. Strange [Car83]. strategy [Bor86]. streamer [HW85]. Street [De 86]. strength [RLH88]. striped [BYNB87, Lan85]. strobe [PCS^{+}85, SHS87]. Strombus [App88]. Strongylocentrotus [SLH87]. structure [WHW88]. structured [Pak89]. Stuart [Moh83]. Student [Mit82b]. Studies [BG85, Joh84, MR85, Mit85, Som89, SJS^{+}89, Thr85, Tro82, Wym86, Bri83, Fin86, God89, HCG86, LT89, SD89, Eve84]. study [BCE86, BK89, Bor86, CC85a, HW85, HOE89, PW86, Kes82d]. sub [BG85]. sub-lethal [BG85]. Subject [Ano82f, Ano85f, Ano86i, Ano87g, Ano88h]. subscription [McI89a]. subsequent [WH89]. subsidiary [DS87]. Subsistence [Wat82]. substrates [DB89]. summer [MWS83]. Sündermann [Lon85]. Suomala [Wym82]. supplement [TNW89]. surface [All89, Fur85a, Fur85b, Ken85, Paw89, Som89, Thr85]. surplus [GSC89]. Surrey [Car83, Car86a, Cus82a, Dah82a, Dah82b, Hop89, Kar84, Kes82a, Kes82d, Kes85b, Lis82, Loc88, Moh83, Nak85, Pee84, Pet86, Ruc86, Smi82, Suo82, Tho88]. survey [EW87]. surveys [BEKV82, Orł89, UB89a]. survival [GOMH85, ST82, ST89, WH89, All83]. Susan [Mit88]. Susanne [Lin87]. Sussex [Fra86]. sustainable [GSC89]. Sutinen [Kes85a]. Suva [Joh84]. swimming [LS89, NC89, WW82]. swordfish [TT89]. Sydney [Thr84]. sympatric [PC89]. Symposium [Joh85, Nau85, Wym83]. system [BEKV82, Wat82]. Systems [Som87].

T [All89, Cus82a, Gjø84a, Kes84d, Pri84b, ZW88]. table [Lan82]. tables [All82a, Car86a, Cha82, Cus82a, Kes82a, Kes84a, Kes84d, Kes85b, Ruc86, Som88]. tagged [DGMO88]. tagging [God89, MYH86]. tags [HW85]. Taivo [Smi82]. Taiwanese [Edw83]. tank [SF85]. target [RLH88]. Tate [GM88, SGB87]. Tavolga [Smi83]. Technical [All89, Fur85a, Fur85b, Mit82a]. technique [Apr87, BCE86]. techniques [PDS86, Mac83, Tho86b]. Technische [Som89]. technology [Da 85a, Da 85b, Kes82a]. telemetry [LS89]. temperate [Lee88]. temperature [DGMO88]. temperatures [LS89, NR87]. Templeton [Pet86]. term [Fag88]. terms [Car83]. Terry [Nau85]. Tertiary [Pol83]. Texas [DGMO88]. text [Fre89]. Their [All83]. theoretical [PW85]. Theory [Gar84, Cus82b, Gar84]. Theragra [ANM87, LG87]. Thermaikos [Pap84]. these [RP82]. Theutenberg [Bor86]. thickness [Fer89]. Thomasson [Kes82d]. those [RP82]. Thunnus [Tho86a]. tidal [Ste88]. tiger [HW85]. Timber [Mit88]. time [BS87, GSC89, God89]. today [Kes82f]. together [Pri84a]. Tokyo [Fin86, Kir86, Mit85, Som86]. tomorrow [Kes82f]. Tongan [Joh84]. Torry [Kes82a]. Tortonese [Daa87a, Daa87b]. towed [Goe82, PM86]. Toxic [UB89b]. toxicology [Wid86]. tracked [Pri84a]. traditional [Tro82]. training [ST89]. transect [Qui85]. transportation [WW82]. trap [BS87]. trapping [JK87]. traps [MR85]. Trawl [EW87, LSC85, Cru86, GM82, GW89, GE86, VWN87, War89]. trawler [Edw83]. trawlers [DS88, WH89]. Trawling [MS85, GM82, NC89, WH89]. trawls [Mac82, SSR89]. trends [BWL87, Fag88, Lis82]. trials [DS89]. triggerfish [MD87]. Trigla [Pap84]. TRIOPS [Som89]. TRIPS [Paw89]. Trivedi [Som88]. Tropical [Paw89, Som89, Bai86, Cus82b, Edw83, Gar84, SJS^{+}89]. Tropicals [Wis84]. trout [All83, MYH86, WB82]. tshawytscha [Fag88, Joh88]. Tsukayama [Som89]. tuna [Tho86a, Dah82b]. twine [Fer89]. two [BC88, CCS86, Cru86, DP89, MS85, McI89a, PC89, RLH88]. two-level [MS85]. Tycooly [Bor86]. typus [DP89].

U [Pee84]. U.K [Gjø84a, Moh83]. U.S [Eve84, Nau85, Pil86, Tho89a, Wis84]. U.S. [Blo87, MD87]. U.S.$3.50 [Joh84]. U.S.A [Gjø84a, Joh85, Mit84, Pol83, Tho86b, Wym82, Wym83, You86]. ultrasonic [LS89]. UNAM [Kes84c]. undersize [BC83]. undersized [MS85]. underwater [GG84]. UNESCO [Daa87b, Daa87a]. Unit [Hee89]. United [Car83, Kes84b]. University [All83, Eve84, Fla82, Joh84, Kes83b, Kes83c, Kes84d, Mac83, Mit82a, Mit84, Mit87, Nau85, Tho89b, Thr84, Wei85]. upwelling [Som89]. urchin [SLH87]. US$0.30 [Tho89a]. US$1.50 [Tho89a]. US$1.80 [Tho89a]. US$10.95 [Tho89a]. US$12.00 [Mit84]. US$15.90 [Kes84b]. US$16.00 [Mac83]. US$17.45 [All83]. US$17.50./£ [Kes83b]. US$28.50 [Gar84]. US$3.00 [Fla82]. US$4.50 [All89]. US$41.00 [Hee89]. US$65.00 [Gjø84b]. US$7.50 [Kes84c]. US$9.00 [McI89a]. US$91.50 [Pri84b]. Use [Apr87, Bai86, BK89, CC85b, God89, Haw82, SGB87]. used [Fer89]. using [GM82, GG84, LS89]. utility [BCE86]. utilization [Jør85, Lis82].

V [Da 85a, Da 85b, Joh85, Mit82a, SY84]. VA [Tho89a]. Val. [AH86]. Valenciennes [DP89]. variability [FS89, You86, Hee89]. Variation [JO87, Jør85, SRRU82, Tan85, War89]. venture [Ken85]. ventures [Mit82a]. Verlag [Fin86, Kes84b, Kir86, Lew85, Lon85, McI89b, Mit85, Nae86, Pri86, Smi83, Som86, Wym86]. Vertebrate [All82a]. vessels [Hop89]. Victoria [SGB87]. vii [Cus82b, Fin86, Mit87]. viral [Eve84]. Virgin [MD87]. Vol [All82a, Cha82, Da 85a, Da 85b, Daa87a, Fin86, Fra86, McI89a, Mit85, Pri86, Wal89, Wym82, Wym86, UB89b]. Volta [Van82]. Volume [Daa87b, Kes85a, Ano88d, Ano89c, McI89a]. Volumen [Kes84c]. VPA [LSM85]. vulgaris [LS89].

W [Cha82, Fra86, Kes83b, Lon85, Mac83, Mit82b, Mit85, Pri89, Smi83, UB89b, Wei85]. WA [Mit84, You86]. Walbaum [BC89]. Walk [Pet86]. walleye [ANM87, BWL87, LG87]. Wanshan [ZW88]. Ward [Gjø84b, Wym83]. Warren [You86]. Washington [All83, Eve84, Fla82, Kes84d, Mit82a, Mit84, Wid86, You86, MWS83]. water [GG84, Mat88, SHS87, WD87]. waters [AH84, AH86, Bai86, Car88, Edw83, GM82, Goe82, HT86, HH88, JO87, Mos88, RP82, All89]. watershed [CCS86]. Weight [Atk89, NNAS89]. weighted [Cri83]. weights [Fag88]. well [Paw89]. Western [Nae86, ANM87, BC83, LSC85]. Westport [Gjø84a, Joh85, Wis84]. Westview [Hee89, Mit82b, Sae82]. whale [Pri89]. whales [Wym84]. which [Fre89]. white [SRRU82, TB88]. whitefish [SN89, SVQ85]. Whitehead [Daa87a, Daa87b]. wildlife [Som87]. Wildwood [Hee89]. Wiley [Fra86, Fre89, Som87]. Wiley-Interscience [Fre89]. William [Kes83b, Som87]. Windsor [Moh83]. wing [PW85, PW86]. Winter [De 86, LS89]. Woodward [Tho89a]. Wooster [You86]. Words [Kes83c]. Worker [Ken85]. Workshop [Gar84, GS87, Kir83]. world [Han82, Mit82a, Car86b].

X [Thr85, Kes85b, Mac83, Mit82b, Pol83, Som87, Tho89a, Wym83]. xi [Car83, Paw89, Wym82]. xii [Lew85, Som89, Wym82]. Xiphias [TT89]. XVI [Smi83]. xx [Mit87, Wei85]. xxxvii [Joh85].

Yami [Dah82b]. year [GM88, You86]. yeild [BH86]. yellow [Pap84]. yellowtail [MD87]. yield [BK89, DS89, GSC89, LG87, Mos88, RL89]. York [Fin86, Kes84b, Kes85a, Kir86, Lew85, Lon85, Mit85, Pol83, Pri84b, Smi83, Som86, Wal89, Wei85, Wym86]. young [Mor86]. Youngs [Kes83b]. Yudanov [Wym82].

zone [Orł89]. Zoophysiology [Pri86]. Zusammenarbeit [Som89].

References

Abu-Hakima:1984:CAR

[AH84] Randa Abu-Hakima. Comparison of aspects of the reproductive biology of Pomadasys, Otolithes and Pampus spp. in Kuwaiti waters. Fisheries Research, 2(3):177–200, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

          sciencedirect.com/science/article/pii/016578368490002X.

Almatar:1986:DAS

[AH86] S. M. Almatar and E. D. Houde. Distribution and abundance of sardine Sardinella fimbriata (Val.) eggs in Kuwait waters of the Arabian Gulf. Fisheries Research, 4(3–4):331–342, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

          1872-6763 (electronic). URL http://www.sciencedirect.com/
          science/article/pii/0165783686900123.

Allen:1981:BRC

[All82a] K. Radway Allen. Book review: Charrs: Salmonid fishes of the genus: Salvelinus. Eugene K. Balon (Editor). Perspectives in Vertebrate Science. Vol. 1. Junk, B.V., The Hague, The Netherlands, 1980. ix + 928 pp., 353 figs., 126 tables, 16 colour plates, cloth, Dfl.400.00/US.$210.60, ISBN 90-6193-701-9. Fisheries Research, 1 (??):187–190, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900230.

Allen:1981:OIP

[All82b] K. Radway Allen. Observations on the identification of population levels giving maximum recruitment. Fisheries Research, 1 (??):251–259, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368190028X.

Allen:1983:BRC

[All83] K. Radway Allen. Book review: The Columbia river salmon and steelhead trout. Their fight for survival: Anthony Netboy, University of Washington Press, Seattle and London 1980, 180 pp., 25 illus., US$17.45, ISBN 0-295-95768-9. Fisheries Research, 2(1): 74–77, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368390108X.

Allen:1989:BRG

[All89] K. Radway Allen. Book review: Growth, mortality and recruitment of commercially important fishes and penaeid shrimps in Indonesian waters: A. Dwiponggo, T. Hariati, S. Banon, M. L. Palomares and D. Pauly, ICLARM Technical Reports 17. Research Institute for Marine Fisheries, Jakarta, Indonesia and the International Center for Living Aquatic Resources Management, 1986, 91 pp., US$4.50 surface, $8.50 airmail, ISBN: 971-10-2226-5; ISSN: 0115-5547. Fisheries Research, 7(1–2):189–191, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900209.

Aneer:1989:HCH

[Ane89] Gunnar Aneer. Herring (Clupea harengus L.) spawning and spawning ground characteristics in the Baltic Sea. Fisheries Research, 8 (2):169–195, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900301.

Alton:1987:CAD

[ANM87] Miles S. Alton, Martin O. Nelson, and Bernard A. Megrey. Changes in the abundance and distribution of walleye pollock (Theragra chalcogramma) in the western Gulf of Alaska (1961–1984). Fisheries Research, 5(2–3):185–197, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900397.

Anonymous:1981:AI

[Ano82a] Anonymous. Author index. Fisheries Research, 1(??):355–357, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900473.

Anonymous:1981:EB

[Ano82b] Anonymous. Editorial Board. Fisheries Research, 1(??):iii, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900011.

Anonymous:1981:Ea

[Ano82c] Anonymous. Erratum. Fisheries Research, 1(??):269, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900345.

Anonymous:1981:Ec

[Ano82d] Anonymous. Erratum. Fisheries Research, 1(??):354, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900461.

Anonymous:1981:P

[Ano82e] Anonymous. Pages 1–361 (1981–1982). Fisheries Research, 1(??): 1–361, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1981:SI

[Ano82f] Anonymous. Subject index. Fisheries Research, 1(??):358–361, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900485.

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. Fisheries Research, 2(1):81, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368390111X.

Anonymous:1983:EB

[Ano83b] Anonymous. Editorial Board. Fisheries Research, 2(1):iii, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368390098X.

Anonymous:1983:PJ

[Ano83c] Anonymous. Pages 1–87 (June 1983). Fisheries Research, 2(1):??, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1983:PS

[Ano83d] Anonymous. Pages 89–166 (September 1983). Fisheries Research, 2 (2):??, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1984:Ab

[Ano84a] Anonymous. Announcement. Fisheries Research, 2(4):325, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900390.

Anonymous:1984:Aa

[Ano84b] Anonymous. Announcements. Fisheries Research, 2(3):240– 242, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900158.

Anonymous:1984:E

[Ano84c] Anonymous. Errata. Fisheries Research, 2(3):239, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783684900146.

Anonymous:1984:PA

[Ano84d] Anonymous. Pages 167–242 (April 1984). Fisheries Research, 2(3): ??, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1984:PS

[Ano84e] Anonymous. Pages 243–325 (September 1984). Fisheries Research, 2(4):??, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1985:AI

[Ano85a] Anonymous. Author index. Fisheries Research, 3(??):387– 389, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368590044X.

Anonymous:1985:EB

[Ano85b] Anonymous. Editorial Board. Fisheries Research, 3(??):iii, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900013.

Anonymous:1985:P

[Ano85c] Anonymous. Pages 1–396 (1985). Fisheries Research, 3(??):1–396, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1985:PNa

[Ano85d] Anonymous. Publisher’s note. Fisheries Research, 3(??):v, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900025.

Anonymous:1985:PNb

[Ano85e] Anonymous. Publisher’s note. Fisheries Research, 3(??):91, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900116.

Anonymous:1985:SI

[Ano85f] Anonymous. Subject index. Fisheries Research, 3(??):390– 396, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900451.

Anonymous:1986:A

[Ano86a] Anonymous. Announcement. Fisheries Research, 4(3–4):366, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900214.

Anonymous:1986:AI

[Ano86b] Anonymous. Author index. Fisheries Research, 4(3–4):367–368, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900226.

Anonymous:1986:BRF

[Ano86c] Anonymous. Book review: Fish migration: Brian A. McKeown. Croom Helm, Beckenham, Kent, England, 1984, 244 pp., £19.95 hardcover, ISBN 0-7099-1761-9. Fisheries Research, 4(2): 169–171, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900421.

Anonymous:1986:EB

[Ano86d] Anonymous. Editorial Board. Fisheries Research, 4(1):iii, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368690024X.

Anonymous:1986:E

[Ano86e] Anonymous. Erratum. Fisheries Research, 4(3–4):365, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900202.

Anonymous:1986:PM

[Ano86f] Anonymous. Pages 1–99 (March 1986). Fisheries Research, 4(1):??, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1986:PJ

[Ano86g] Anonymous. Pages 101–180 (July 1986). Fisheries Research, 4(2): ??, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1986:PD

[Ano86h] Anonymous. Pages 181–375 (December 1986). Fisheries Research, 4(3–4):??, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1986:SI

[Ano86i] Anonymous. Subject index. Fisheries Research, 4(3–4):369–375, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900238.

Anonymous:1987:AI

[Ano87a] Anonymous. Author index. Fisheries Research, 5(4):421–423, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900609.

Anonymous:1987:EBa

[Ano87b] Anonymous. Editorial Board. Fisheries Research, 5(1):iii, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900117.

Anonymous:1987:EBb

[Ano87c] Anonymous. Editorial Board. Fisheries Research, 6(1):iii, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900014.

Anonymous:1987:PM

[Ano87d] Anonymous. Pages 1–103 (May 1987). Fisheries Research, 5(1):??, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1987:PN

[Ano87e] Anonymous. Pages 1–103 (November 1987). Fisheries Research, 6 (1):??, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1987:PA

[Ano87f] Anonymous. Pages 331–428 (August 1987). Fisheries Research, 5 (4):??, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1987:SI

[Ano87g] Anonymous. Subject index. Fisheries Research, 5(4):424–428, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900610.

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. Fisheries Research, 6(3):304, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783688900239.

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcement. Fisheries Research, 6(4):399, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783688900100.

Anonymous:1988:Ac

[Ano88c] Anonymous. Authorindex. Fisheries Research, 6(4):401–403, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900112.

Anonymous:1988:CFR

[Ano88d] Anonymous. Contents of fisheries research, volume 6. Fisheries Research, 6(4):413–415, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900136.

Anonymous:1988:PJ

[Ano88e] Anonymous. Pages 105–199 (January 1988). Fisheries Research, 6 (2):??, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1988:PM

[Ano88f] Anonymous. Pages 201–304 (May 1988). Fisheries Research, 6(3): ??, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1988:PS

[Ano88g] Anonymous. Pages 305–415 (September 1988). Fisheries Research, 6(4):??, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1988:SI

[Ano88h] Anonymous. Subject index. Fisheries Research, 6(4):405–411, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900124.

Anonymous:1989:A

[Ano89a] Anonymous. Announcement. Fisheries Research, 8(2):203, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900349.

Anonymous:1989:AI

[Ano89b] Anonymous. Author index. Fisheries Research, 7(4):389–391, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900726.

Anonymous:1989:CFR

[Ano89c] Anonymous. Contents of fisheries research, volume 7. Fisheries Research, 7(4):393–395, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900738.

Anonymous:1989:Ca

[Ano89d] Anonymous. Corrigendum. Fisheries Research, 7(1–2):193, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900222.

Anonymous:1989:Cb

[Ano89e] Anonymous. Corrigendum. Fisheries Research, 7(3):299, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900635.

Anonymous:1989:EBa

[Ano89f] Anonymous. Editorial Board. Fisheries Research, 7(1–2):iii, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900015.

Anonymous:1989:EBb

[Ano89g] Anonymous. Editorial Board. Fisheries Research, 8(1):iii, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900350.

Anonymous:1989:E

[Ano89h] Anonymous. Erratum. Fisheries Research, 7(1–2):194, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900234.

Anonymous:1989:PF

[Ano89i] Anonymous. Pages 1–200 (February 1989). Fisheries Research, 7 (1–2):??, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1989:PJb

[Ano89j] Anonymous. Pages 1–91 (July 1989). Fisheries Research, 8(1):??, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1989:PM

[Ano89k] Anonymous. Pages 201–299 (May 1989). Fisheries Research, 7(3): ??, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1989:PJa

[Ano89l] Anonymous. Pages 301–395 (June 1989). Fisheries Research, 7(4): ??, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1989:PN

[Ano89m] Anonymous. Pages 93–203 (November 1989). Fisheries Research, 8 (2):??, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Angelsen:1987:IFD

[AO87] Karl Kr. Angelsen and Steinar Olsen. Impact of fish density and effort level on catching efficiency of fishing gear. Fisheries Research, 5(2–3):271–278, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900464.

Amarasinghe:1986:AFE

[AP86] U. S. Amarasinghe and T. J. Pitcher. Assessment of fishing effort in Parakrama Samudra, an ancient man-made lake in Sri Lanka. Fisheries Research, 4(3–4):271–282, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900081.

Appeldoorn:1988:ADG

[App88] Richard S. Appeldoorn. Age determination, growth, mortality and age of first reproduction in adult Queen conch, Strombus gigas L., off Puerto Rico. Fisheries Research, 6(4):363–378, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900057.

Aprahamian:1987:UBT

[Apr87] Miran W. Aprahamian. Use of the burning technique for age determination in eels (Anguilla anguilla (L.)) derived from the stocking of elvers. Fisheries Research, 6(1):93–96, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900099.

Atkinson:1989:WLR

[Atk89] D. B. Atkinson. Weight–length relationships of roundnose grenadier (Coryphaenoides rupestris Gunn.) in different areas of the North Atlantic. Fisheries Research, 7(1–2):65–72, February 1989. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783689900076.

Bailey:1981:DCS

[BA82] Kevin M. Bailey and David G. Ainley. The dynamics of California sea lion predation on Pacific hake. Fisheries Research, 1 (??):163–176, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900187.

Bailey:1986:BRF

[Bai86] R. S. Bailey. Book review: Fish population dynamics in tropical waters: a manual for use with programmable calculators: D. Pauly. International Center for Living Aquatic Resources Management (I.C.L.A.R.M.), M.C.C., P.O. Box 1501, Makati, Manila, Philippines, 1984, 325 pp., hardback, illustrated, price not given, ISBN 971-10-2203-6. Fisheries Research, 4(2):171– 173, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900445.

Brown:1983:FAR

[BC83] R. S. Brown and N. Caputi. Factors affecting the recapture of undersize western rock lobster Panulirus cygnus George returned by fishermen to the sea. Fisheries Research, 2(2):103– 128, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683901133.

Boy:1988:SDG

[BC88] Vincent Boy and Alain J. Crivelli. Simultaneous determination of gillnet selectivity and population age-class distribution for two cyprinids. Fisheries Research, 6(4):337–345, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783688900033.

Bowering:1989:DRA

[BC89] W. R. Bowering and A. K. Chumakov. Distribution and relative abundance of Greenland halibut (Reinhardtius hippoglossoides (Walbaum)) in the Canadian Northwest Atlantic from Davis Strait to the northern Grand Bank. Fisheries Research, 7(4): 301–327, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900647.

Beninger:1986:UAC

[BCE86] Peter G. Beninger, Lorette Chiasson, and Robert W. Elner. The utility of artificial collectors as a technique to study benthic settlement and early juvenile growth of the rock crab, Cancer irroratus. Fisheries Research, 4(3–4):317–329, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900111.

Bayer:1987:SEO

[BD87] Robert C. Bayer and Peter C. Daniel. Safety and efficacy of oxytetracycline for control of gaffkemia in the American lobster (Homarus americanus). Fisheries Research, 5(1):71–81, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900166.

Blindheim:1981:SBD

[BEKV82] Johan Blindheim, Per K. Eide, Hans P. Knudsen, and Gudmund Vestnes. A ship-borne data logging and processing system for acoustic fish surveys. Fisheries Research, 1(??):141–153, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900163.

Bernard:1988:PFG

[Ber88a] F. R. Bernard. Potential fishery for the gooseneck barnacle Pollicipes polymerus (Sowerby, 1833) in British Columbia. Fisheries Research, 6(3):287–298, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900203.

Bertrand:1988:SHH

[Ber88b] Jacques Bertrand. Selectivity of hooks in the handline fishery of the Saya de Malha banks (Indian Ocean). Fisheries Research, 6(3): 249–255, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900173.

Bhaskar:1985:SFL

[BG85] M. Bhaskar and S. Govindappa. Studies on fish liver carbohydrate metabolism during acclimation to sub-lethal acidic and alkaline media. Fisheries Research, 3(??):343–350, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900323.

Botsford:1986:SOY

[BH86] Louis W. Botsford and Roderick C. Hobbs. Static optimization of yeild per recruit with reproduction and fishing costs. Fisheries Research, 4(3–4):181–189, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900019.

Bergstad:1987:LHE

[BJD87] O. A. Bergstad, T. Jørgensen, and O. Dragesund. Life history and ecology of the gadoid resources of the Barents Sea. Fisheries Research, 5(2–3):119–161, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900373.

Berkes:1989:CSY

[BK89] Fikret Berkes and Mina Kislalioglu. A comparative study of yield, investment and energy use in small-scale fisheries: Some considerations for resource planning. Fisheries Research, 7(3): 207–224, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900556.

Blay:1985:OBF

[Bla85] John Blay. Observations on the balance in fish populations in a small reservoir in Ghana. Fisheries Research, 3(??):1– 11, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900037.

Blomo:1987:DEI

[Blo87] Vito J. Blomo. Distribution of economic impacts from proposed conservation measures in the U.S. Atlantic menhaden fishery. Fisheries Research, 5(1):23–38, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900130.

Borgese:1986:BRE

[Bor86] E. Mann Borgese. Book review: The evolution of the law of the sea. A study of resources and strategy, with special regard to the polar areas: by Bo Johnson Theutenberg. Tycooly International Publishing, Dublin, 1984, 261 pp., price hardcover £16.50/US $25.00, softcover £8.50/US $12.50, ISBN hardcover 0-86346-041-0, softcover 0-86346-042-9. Fisheries Research, 4(1): 82–84, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900330.

Borgstrom:1989:DEG

[Bor89] Reidar Borgström. Direct estimation of gill-net selectivity for roach (Rutilus rutilus (L.)) in a small lake. Fisheries Research, 7(3): 289–298, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900623.

Briggs:1983:NSS

[Bri83] R. P. Briggs. Net selectivity studies in the Northern Ireland Nephrops fishery. Fisheries Research, 2(1):29–46, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783683901017.

Briggs:1985:CCN

[Bri85] R. P. Briggs. Catch composition in the Northern Ireland Nephrops fishery. Fisheries Research, 3(??):47–60, ???? 1985. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783685900074.

Briggs:1986:GRM

[Bri86] R. P. Briggs. A general review of mesh selection for Nephrops norvegicus (L.). Fisheries Research, 4(1):59–73, March 1986. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783686900287.

Brown:1983:BRN

[Bro83] Bradford E. Brown. Book review: Nova Scotia Fisheries Atlas: Maritime Resource Management Service (MRMS), Nova Scotia, Canada, 1982. 44 pp., Can. $12.50, No ISBN. Fisheries Research, 2(2):162–163, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901182.

Boutillier:1987:ETD

[BS87] J. A. Boutillier and N. A. Sloan. Effect of trap design and soak time on catches of the British Columbia prawn (Pandalus platyceros). Fisheries Research, 6(1):69–79, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900075.

Bakkala:1987:HTA

[BWL87] Richard G. Bakkala, Vidar G. Wespestad, and Loh-Lee Low. Historical trends in abundance and current condition of walleye pollock in the eastern Bering Sea. Fisheries Research, 5(2–3):199– 215, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900403.

Brown:1987:PRP

[BYNB87] Christopher L. Brown, Graham Young, Richard S. Nishioka, and Howard A. Bern. Preliminary report on the physiological status of striped bass in the Carquinez Strait die-off. Fisheries Research, 6(1):5–16, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900038.

Carrothers:1983:BRG

[Car83] P. J. G. Carrothers. Book review: Glossary of United Kingdom fishing gear terms: J. P. Bridger, J. J. Foster, A. R. Margetts and E. S. Strange. Fishing News Books, Surrey, 1981, xi + 114 pp., 33 illus., £15, + £1.50 post/packing, ISBN 0-85238-119-0. Fisheries Research, 2(1):66–70, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901054.

Carrothers:1986:BRFa

[Car86a] P. J. G. Carrothers. Book review: Fibre ropes for fishing gear: (FAO Fishing Manuals) by Gerhard Klust. Fishing News Books, Farnham, Surrey, Gt. Britain, 1983, 199 pp., 62 illustrations, 41 tables, 252 references, price £9.50, ISBN-0-85238-123-9. Fisheries Research, 4(1):85–93, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900354.

Carrothers:1986:BRFb

[Car86b] P. J. G. Carrothers. Book review: Fish catching methods of the world: A. von Brandt. 3rd Edn. Fishing News Books, Farnham, 1984, 418 pp., 733 figs., 690 refs., £27.50, ISBN 0-85238-125-5. Fisheries Research, 4(2):176–180, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900470.

Caramelo:1988:ECM

[Car88] Ana Maria Caramelo. Effects of a change in mesh size on Nephrops stocks in Portuguese waters. Fisheries Research, 6(4):379–391, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900069.

Carvajal:1985:SAI

[CC85a] Juan Carvajal and Pedro E. Cattan. A study of the anisakid infection in the Chilean hake, Merluccius gayi (Guichenot, 1848). Fisheries Research, 3(??):245–250, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900256.

Chien:1985:MDM

[CC85b] Yew-Hu Chien and Richard E. Condrey. A modification of the DeLury method for use when natural mortality is not negligible. Fisheries Research, 3(??):23–28, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900050.

Couturier:1986:ISA

[CCS86] C. Y. Couturier, L. Clarke, and A. M. Sutterlin. Identification of spawning areas of two forms of Atlantic salmon (Salmo salar L.) inhabiting the same watershed. Fisheries Research, 4(2): 131–144, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368690038X.

Chapman:1981:BRM

[Cha82] Donald Chapman. Book review: Man and fisheries on an Amazon frontier: Michael Goulding. W. Junk BV, The Hague. Developments in Hydrobiology, Vol. 4, 1981. 137 pp., 96 figs., 5 tables, Dfl. 90.00/US $47.50, ISBN 90-6193-755-8. Fisheries Research, 1 (??):265–266, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368190031X.

Crittenden:1983:ELD

[Cri83] Robert N. Crittenden. An evaluation of the Leslie–DeLury method and of a weighted method for estimating the size of a closed population. Fisheries Research, 2(2):149–158, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783683901157.

Crozier:1987:OCC

[Cro87] W. W. Crozier. Occurrence of Contracaecum clavatum Rudolphi in angler-fish (Lophius piscatorius L.) from the North Irish Sea. Fisheries Research, 5(1):83–90, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900178.

Crozier:1989:AGA

[Cro89] W. W. Crozier. Age and growth of angler-fish (Lophius piscatorius L.) in the north Irish Sea. Fisheries Research, 7(3): 267–278, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368990060X.

Crutchfield:1986:PCS

[Cru86] Stephen R. Crutchfield. Personal computer simulations of two New England trawl fisheries. Fisheries Research, 4(2):157–165, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900408.

Cushing:1981:BRM

[Cus82a] D. H. Cushing. Book review: Marine fisheries ecosystem, its quantitative evaluation and management: T. Laevastu and H. A. Larkins. Fishing News Books, Farnham, Surrey, Great Britain, 1981. 155 pp., 53 illus., 18 tables, 246 × 189 mm. £15.50, ISBN 0-85238-116-6. Fisheries Research, 1(??):262–265, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900308.

Cushing:1981:BRT

[Cus82b] D. H. Cushing. Book review: Theory and management of tropical multispecies stocks: Daniel Pauly. ICLARM, Metro Manila, The Philippines, 1979, paperback, vii + 35 pp., ISSN 0115-4389. Fisheries Research, 1(??):182–184, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900217.

DaCasta:1985:BRH

[Da 85a] Antonio Da Casta. Book review: Handbook of fishery technology: Vol. 4. V. M. Novikov (Editor). A. A. Balkema Publishers, Rotterdam, The Netherlands, 1984, 490 pp., cloth Dfl. 75.00, US $28.00, £17.50, ISBN 90-6191-421-3. Fisheries Research, 3(??): 375–376, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900360.

DaCosta:1985:BRH

[Da 85b] Antonio Da Costa. Book review: Handbook of fishery technology: Vol. 1. V. M. Novikov (Editor). A. A. Balkema Publishers, Rotterdam, The Netherlands, 1983, 504 pp., cloth Dfl. 75.00, US $28.00, £17.50, ISBN 90-6191-409-4. Fisheries Research, 3(??): 375–376, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900359.

Daan:1987:BRFa

[Daa87a] Niels Daan. Book review: Fishes of the North-eastern Atlantic and the Mediterranean: Vol. 1. P. J. P. Whitehead, M.-L. Bauchot, J.-C. Hureau, J. Nielsen and E. Tortonese (Editors). The Unesco Press, 1984, 510 pp., FF250, £37.50, $63.50, ISBN 92-3-002215-2. Fisheries Research, 5(4):416–417, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900580.

Daan:1987:BRFb

[Daa87b] Niels Daan. Book review: Fishes of the North-eastern Atlantic and the Mediterranean: Volume II. P. J. P. Whitehead, M.-L. Bauchot, J.-C. Hureau, J. Nielsen, E. Tortonese (Editors). UNESCO Press, Paris, 1986, pp. 510–1007, FF 250, Dfl 105.00, ISBN 92-3-002308-6. Fisheries Research, 6(1):97, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900105.

Dahle:1981:BRF

[Dah82a] E. A. Dahle. Book review: FAO investigates ferro-cement fishing craft: John Fyson (Editor). Fishing News Books Ltd., Surrey, Great Britain, 1980. 200 pp., 158 illus., 276 × 219 mm, hardback, £15.50, ISBN 0-85238-061-5. Fisheries Research, 1(??):75, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900114.

Dahle:1981:BRT

[Dah82b] E. A. Dahle. Book review: Tuna fishing with pole and line: M. Ben-Yami (Editor). Published by Fishing News Books Ltd., Farnham, Surrey, Great Britain, by arrangement with FAO. 168 pp., 129 illus., 234 × 156 mm, £6.50, ISBN 0-85238-111-5. Fisheries Research, 1(??):352, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900448.

Davis:1986:MPB

[Dav86] T. L. O. Davis. Migration patterns in barramundi, Lates calcarifer (Bloch), in Van Diemen Gulf, Australia, with estimates of fishing mortality in specific areas. Fisheries Research, 4(3–4):243– 258, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900068.

Daniel:1989:FBC

[DB89] Peter C. Daniel and Robert C. Bayer. Fish byproducts as chemoattractant substrates for the American lobster (Homarus americanus): Concentration, quality and release characteristics. Fisheries Research, 7(4):367–383, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900684.

DeGroot:1986:BRL

[De 86] S. J. De Groot. Book review: The Lammas Drave and the Winter Herrin’. A history of the herring fishing from east fife: Peter Smith. John Donald Publishers, 138 St. Stephen Street, Edingburgh EH3 5AA, Gt. Britain, 1985, 166 pp., illustrated, price £5.95, ISBN 0-85976-125-8. Fisheries Research, 4(2):173– 174, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900457.

Dayaratne:1986:AGF

[DG86] Pauline Dayaratne and Jakob Gjøsaeter. Age and growth of four Sardinella species from Sri Lanka. Fisheries Research, 4(1):1– 33, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900251.

Doerzbacher:1988:TCB

[DGMO88] J. F. Doerzbacher, A. W. Green, G. C. Matlock, and H. R. Osburn. A temperature compensated von Bertalanffy growth model for tagged red drum and black drum in Texas bays. Fisheries Research, 6(2):135–152, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783688900331.

Dickson:1989:CGE

[Dic89a] William Dickson. Cod gillnet effectiveness related to local abundance, availability and fish movement. Fisheries Research, 7(1– 2):127–148, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368990012X.

Dickson:1989:CGS

[Dic89b] William Dickson. Cod gillnet simulation model. Fisheries Research, 7(1–2):149–174, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900131.

Dunn:1987:REL

[DM87] Jean R. Dunn and Ann C. Matarese. A review of the early life history of Northeast Pacific gadoid fishes. Fisheries Research, 5 (2–3):163–184, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900385.

Djama:1989:CSA

[DP89] T. Djama and T. J. Pitcher. Comparative stock assessment of two sciaenid species, Pseudotolithus typus (Bleeker, 1863) and Pseudotolithus senegalensis (Valenciennes, 1833) off Cameroon. Fisheries Research, 7(1–2):111–125, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900118.

DeSilva:1987:NFR

[DS87] Sena S. De Silva and H. K. G. Sirisena. New fish resources of reservoirs in Sri Lanka: Feasibility of introduction of a subsidiary gillnet fishery for minor cyprinids. Fisheries Research, 6(1):17– 34, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368790004X.

Devaraj:1988:EPM

[DS88] M. Devaraj and Paralkar Smita. Economic performance of mechanised trawlers in the state of Kerala, India. Fisheries Research, 6 (3):271–286, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900197.

DeSilva:1989:NFR

[DS89] Sena S. De Silva and H. K. G. Sirisena. New fish resources of reservoirs in Sri Lanka. 3. Results of commercial-scale trials and yield estimates of a gill-net fishery for minor cyprinids. Fisheries Research, 7(3):279–287, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900611.

Edwards:1983:TPT

[Edw83] R. R. C. Edwards. The Taiwanese pair trawler fishery in tropical Australian waters. Fisheries Research, 2(1):47–60, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783683901029.

Erkoyuncu:1989:EAS

[EO89] I. Erkoyuncu and E. Ozdamar. Estimation of the age, size and sex composition and growth parameters of anchovy, Engraulis encrasicolus (L.) in the Black Sea. Fisheries Research, 7(3):241– 247, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900581.

Evelyn:1984:BRB

[Eve84] T. P. T. Evelyn. Book review: Bacterial and viral diseases of fish. Molecular studies: Jorge H. Crosa (Editor), A Washington Sea Grant Publication, University of Washington, Seattle, 1983, 86 pp., illustrated, U.S. $5.00. Fisheries Research, 2(4):316–318, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900341.

Engaas:1987:TPD

[EW87] A. Engås and C. W. West. Trawl performance during the Barents Sea cod and haddock survey: potential sources of gearrelated sampling bias. Fisheries Research, 5(2–3):279–286, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900476.

Fagen:1988:LTT

[Fag88] Robert Fagen. Long-term trends in maximum size of sportcaught Chinook salmon (Oncorhynchus tshawytscha): a dataanalytic approach to weights of first-prize fish in four Southeastern Alaska salmon derbies. Fisheries Research, 6(2):125– 134, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368890032X.

Ferro:1989:OMT

[Fer89] R. S. T. Ferro. Objective measurement of the thickness of netting twine used in the fishing industry. Fisheries Research, 8 (2):103–112, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900258.

Ferno:1983:EEB

[FH83] Anders Fernö and Ingvar Huse. The effect of experience on the behaviour of cod (Gadus morhua L.) towards a baited hook. Fisheries Research, 2(1):19–28, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901005.

Finch:1986:BRF

[Fin86] Roland Finch. Book review: Fishery management, Lecture notes on coastal and estuarine studies. Vol. 10, by J. L. McHugh. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1984, vii + 207 pp., soft cover, price DM 43/US$15.10, ISBN 978-0-387-96062-3. Fisheries Research, 4(2):167–168, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578368690041X.

Flanagan:1981:BRE

[Fla82] Brian P. Flanagan. Book review: Extended national fisheries jurisdiction (palliative or panacea?): Roy I. Jackson. University of Washington Press, 1982, 30 pp., paperback, US$3.00. Fisheries Research, 1(??):353–354, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368190045X.

Foote:1981:EAE

[Foo82] Kenneth G. Foote. Energy in acoustic echoes from fish aggregations. Fisheries Research, 1(??):129–140, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900151.

Francis:1986:BRF

[Fra86] Robert C. Francis. Book review: Fish stock assessment, a manual of basic methods: by J. A. Gulland. FAO/Wiley Series on Food and Agriculture, Vol. 1. John Wiley, Chichester, W. Sussex, Gt. Britain, 1983, approx. 224 pp., approx. US $31.90/£15.95, ISBN 0-471-90027-3. Fisheries Research, 4(1):76– 78, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900305.

French:1989:BRD

[Fre89] E. L. French. Book review: Dictionary of microbiology and molecular biology: Second Edition. P. Singleton and D. Sainsbury, A Wiley-Interscience Publication, Wiley, Chichester, 1987, 972 pp. of text, 46 pp. of appendixes which include 22 detailed fermentation or biosynthetic pathways, £69.00, ISBN: 0-0471-91114-3. Fisheries Research, 7(1–2):188–189, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900192.

Ferro:1987:DNC

[FS87] R. S. T. Ferro and P. A. M. Stewart. The drag of nylon codends. Fisheries Research, 5(4):331–347, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900518.

Fagen:1989:HLC

[FS89] Robert Fagen and William W. Smoker. How large-capacity hatcheries can alter interannual variability of salmon production. Fisheries Research, 8(1):1–11, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900362.

Furtado:1985:BRSa

[Fur85a] Jose I. Furtado. Book review: Small-scale fisheries of San Miguel Bay, Philippines: Biology and stock assessment: Daniel Pauly and Antonio N. Mines (Editors). ICLARM Technical Report No. 7, ICLARM, Metro Manila, Philippines, 1982, 124 pp., price US $9.00 surface, US $16.00 airmail, ISSN 0115-5547. Fisheries Research, 3(??):379–380, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783685900396.

Furtado:1985:BRSb

[Fur85b] Jose I. Furtado. Book review: Small-scale fisheries of San Miguel Bay, Philippines: Occupational and geographic mobility: by C. Bailey, ICLARM, Metro Manila, Philippines, ICLARM Technical Report No. 10, 1982, 57 pp., price US $7.00 surface, US $11.50 airmail, ISSN 0115-5547. Fisheries Research, 3(??):379– 380, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900402.

Garcia:1984:BRT

[Gar84] S. Garcia. Book review: Theory and management of tropical fishery: D. Pauly and G. Murphy (Editors). Proceedings of the ICLARM — CSIRO Workshop on the Theory and Management of Tropical Multispecies Stocks, 12–21 January 1981, Cronulla, Australia, published by ICLARM, Manila, Philippines, 1982, 360 pp., paperback price US$28.50, ISBN 971-0400-22-3. Fisheries Research, 2(3): 229–232, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900109.

Gunderson:1986:DPS

[GE86] Donald R. Gunderson and Ian E. Ellis. Development of a plumb staff beam trawl for sampling demersal fauna. Fisheries Research, 4 (1):35–41, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900263.

Gilat:1984:SFO

[GG84] E. Gilat and A. Gelman. On the sharks and fishes observed using underwater photography during a deep-water cruise in the eastern Mediterranean. Fisheries Research, 2(4):257–271, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900298.

Gjosaeter:1984:BRF

[Gjø84a] Jakob Gjøsæter. Book review: Fisheries Ecology: by T. J. Pitcher and P. J. Hart. Published in America by AVI Publishing, Westport, CT, published in Europe by Croom Helm, Kent, U.K., 1982, 414 pp., price US $32.50, £16.95, ISBN 0-87055-405-0. (U.S.A.), ISBN 0-85664-894-9. (U.K.). Fisheries Research, 2(4):315–316, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368490033X.

Gjosaeter:1984:BRP

[Gjø84b] Jakob Gjøsaeter. Book review: Predators and Prey in Fishes: By David I. G. Noakes, David G. Lindquist, Gene S. Helfman and J. A. Ward. Developments in Environmental Biology of Fishes 2. Junk, The Hague, The Netherlands, 1983, 230 pp., price Dfl. 150.00 (approx. US$65.00), ISBN 90-6193-822-8. Fisheries Research, 2(3): 217–218, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900043.

Gibbs:1981:AET

[GM82] P. J. Gibbs and J. Matthews. Analysis of experimental trawling using a miniature otter trawl to sample demersal fish in shallow estuarine waters. Fisheries Research, 1(??):235–249, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900278.

Godo:1987:GMN

[GM87] Olav Rune Godø and Erlend Moksness. Growth and maturation of Norwegian coastal cod and Northeast Arctic cod under different conditions. Fisheries Research, 5(2–3):235–242, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900439.

Gwyther:1988:GRN

[GM88] D. Gwyther and P. E. McShane. Growth rate and natural mortality of the scallop Pecten alba Tate in Port Phillip Bay, Australia, and evidence for changes in growth rate after a 20-year period. Fisheries Research, 6(4):347–361, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900045.

Godo:1989:UTS

[God89] Olav Rune Godø. The use of tagging studies to determine the optimal time for estimating acoustic abundance of spawning cod. Fisheries Research, 8(2):129–140, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900271.

Goeden:1981:TIP

[Goe82] Gerald B. Goeden. A towed instrument package for fisheries research in great barrier reef waters. Fisheries Research, 1(??):35– 44, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900059.

Green:1985:ESR

[GOMH85] A. W. Green, H. R. Osburn, G. C. Matlock, and H. E. Hegen. Estimated survival rates for immature red drum in northwest Gulf of Mexico bays. Fisheries Research, 3(??):263–277, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368590027X.

Gee:1981:RSC

[GR82] Alun S. Gee and Philip J. Radford. The regulation of stock characteristics in a simulated Atlantic salmon population. Fisheries Research, 1(??):105–116, ???? 1981–1982. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900138.

Gunderson:1987:WCB

[GS87] Donald R. Gunderson and Svein Sundby. Workshop on comparative biology, assessment, and management of Gadoids from the North Pacific and Atlantic oceans. Fisheries Research, 5(2–3): 105–110, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368790035X.

Garcia:1989:ESP

[GSC89] S. Garcia, P. Sparre, and J. Csirke. Estimating surplus production and maximum sustainable yield from biomass data when catch and effort time series are not available. Fisheries Research, 8(1): 13–23, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900374.

Glass:1989:CRF

[GW89] C. W. Glass and C. S. Wardle. Comparison of the reactions of fish to a trawl gear, at high and low light intensities. Fisheries Research, 7(3):249–266, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900593.

Hall:1984:MRM

[Hal84] Steven Hall. A multiple regression model of oyster growth. Fisheries Research, 2(3):167–175, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783684900018.

Hansen:1981:MWF

[Han82] Terje Hansen. A model of the world fish meal market. Fisheries Research, 1(??):271–288, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900357.

Hannesson:1987:OCC

[Han87] Rögnvaldur Hannesson. Optimal catch capacity and fishing effort in deterministic and stochastic fishery models. Fisheries Research, 5(1):1–21, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900129.

Hawkins:1981:BRF

[Haw82] A. D. Hawkins. Book review: Fish behaviour and its use in the capture and culture of fishes: J. E. Bardach, J. J. Magnuson, R. C. May and J. M. Reinhart (Editors). International Center for Living Aquatic Resources Management, Manila, Philippines, 1980. ICLARM Conference Proceedings No. 5, 512 pp., paperback, ISBN 971-02-0003-8. Price US $15.00. Fisheries Research, 1(??): 267–269, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900333.

Hay:1986:ESP

[HCG86] D. E. Hay, K. D. Cooke, and C. V. Gissing. Experimental studies of Pacific herring gillnets. Fisheries Research, 4(3–4):191– 211, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900020.

Heessen:1989:BRV

[Hee89] Henk Heessen. Book review: Variability and management of large marine ecosystems: K. Sherman and L. M. Alexander (Editors) Westview Press, Distributed by Wildwood Distribution Service, Unit 3, Lower Farnham Road, Aldershot, Hants GU12 4DY, Gt. Britain, 1986, 319 pp., US$41.00, ISBN: 0-8133-7145-7. Fisheries Research, 7(1–2):184–185, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900167.

Hanumara:1987:ECF

[HH87] R. Choudary Hanumara and Nadine A. Hoenig. An empirical comparison of a fit of linear and non-linear models for seasonal growth in fish. Fisheries Research, 5(4):359–381, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900531.

HjaltiiJakupsstovu:1988:GSM

[HH88] Stein HjaltiíJákupsstovu and Tore Haug. Growth, sexual maturation, and spawning season of Atlantic halibut, Hippoglossus hippoglossus, in Faroese waters. Fisheries Research, 6(3):201– 215, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900148.

Hockin:1989:RSM

[HOE89] D. C. Hockin, K. O’Hara, and J. W. Eaton. A radiotelemetric study of the movements of grass carp in a British canal. Fisheries Research, 7(1–2):73–84, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900088.

Hopper:1989:BRD

[Hop89] A. G. Hopper. Book review: Design of small fishing vessels: J. Fyson (Editor), Fishing News Books Ltd, Farnham, Surrey, 1987, 320 pp., £28.50, ISBN:0-85238-133-6. Fisheries Research, 7(1–2): 192, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900210.

Haug:1986:CSA

[HT86] Tore Haug and Jonas Tjemsland. Changes in size- and agedistributions and age at sexual maturity in Atlantic halibut, Hippoglossus hippoglossus, caught in North Norwegian waters. Fisheries Research, 4(2):145–155, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900391.

Hill:1985:LSE

[HW85] B. J. Hill and T. J. Wassenberg. A laboratory study of the effect of streamer tags on mortality, growth, moulting and duration of nocturnal emergence of the tiger prawn Penaeus esculentus (Haswell). Fisheries Research, 3(??):223–235, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900232.

Jaffar:1989:ACE

[JA89] M. Jaffar and M. Ashraf. Arsenic concentrations in the edible muscle of commercial marine fish species from the Arabian Sea. Fisheries Research, 7(1–2):11–16, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900039.

Jakobsen:1987:CCN

[Jak87] Tore Jakobsen. Coastal cod in Northern Norway. Fisheries Research, 5(2–3):223–234, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900427.

Jensen:1981:ACC

[Jen82] A. L. Jensen. Adjusting catch curves for gill net selection with the logistic distribution. Fisheries Research, 1(??):155–162, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900175.

Jacobsen:1987:SSB

[JK87] Terri Jacobsen and James A. Kushlan. Sources of sampling bias in enclosure fish trapping: Effects on estimates of density and diversity. Fisheries Research, 5(4):401–412, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900555.

Jurvelius:1984:NPF

[JLL84] Juha Jurvelius, Torfinn Lindem, and Jarmo Louhimo. The number of pelagic fish in Lake Paasivesi, Finland, monitored by hydroacoustic methods. Fisheries Research, 2(4):273–283, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900304.

Jakobsen:1987:VRM

[JO87] Tore Jakobsen and Steinar Olsen. Variation in rates of migration of saithe from Norwegian waters to Iceland and Faroe Islands. Fisheries Research, 5(2–3):217–222, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900415.

Johannes:1984:BRT

[Joh84] R. E. Johannes. Book review: The Tongan fishermen: By S. Halapua. Published in 1982 by the Institute of Pacific Studies in association with the Institute of Marine Resources, University of the South Pacific. Available for 3 Fijian dollars (about U.S.$3.50) plus postage from the Institute of Pacific Studies, University of the South Pacific, Suva, Fiji. Fisheries Research, 2(3):

224–226, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900079.

Johnson:1985:BRF

[Joh85] H. T. Johnson. Book review: Freshwater crayfish: V. Charles R. Goldman (Editor). Papers from the Fifth International Symposium on Freshwater Crayfish, Davis, California, U.S.A., 1981, AVI Publishing Company, Westport, CT, 1983, xxxvii + 569 pp., hardback $35.00, ISBN 0-87055-438-7. Fisheries Research, 3(??): 178–181, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900207.

Johnson:1988:ENN

[Joh88] Steven L. Johnson. The effects of the 1983 El Niño on Oregon’s Coho (Oncorhynchus kisutch) and Chinook (O. tshawytscha) salmon. Fisheries Research, 6(2):105–123, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900318.

Jorgensen:1985:CNU

[Jør85] Leif Jørgensen. Carbon and nitrogen utilization in developing eggs and larvae of cod (Gadus morhua, L.) and variation among different parents. Fisheries Research, 3(??):337– 342, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900311.

Karlsen:1984:BRN

[Kar84] Ludvig Karlsen. Book review: Netting materials for fishing gear: 2nd edn. Gerhard Klust. An FAO Fishing Manual. Fishing News Books Ltd., Farnham, Surrey, Gt. Britain, 1982, 192 pp., 61 illustrations, £8.00 plus 80 p post/packing. ISBN: 0-85238-118-2. Fisheries Research, 2(3):228–229, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783684900092.

Kent:1985:BRJ

[Ken85] George Kent. Book review: A Japanese fishing joint venture: Worker experience and national development in the Solomon Islands: by Sarah K. Meltzoff and Edward S. LiPuma. Manila, International Center for Living Aquatic Resources Management, 1983, 63 pp., price US $6.55 surface, US $10.90 airmail, ISSN 0115-5547. Fisheries Research, 3(??):382–383, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900426.

Kesteven:1981:BRA

[Kes82a] G. L. Kesteven. Book review: Advances in fish science and technology: J. J. Connell (Director) and staff of Torry Research Station (Editors). Fishing News Books Ltd. Surrey, Great Britain, 1980. 528 pp., 400 illus., 279 tables, £39.50, ISBN 0-85238-108-5. Fisheries Research, 1(??):71–72, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900084.

Kesteven:1981:BRD

[Kes82b] G. L. Kesteven. Book review: Derecho pesquero: Academia International de Derecho Pesquero. No. 1. Lic. Carlos J. Sierra, Apdo. Postal 49-004, Mexico, 06060 D. F. Mexico un. Ano $600.00. Fisheries Research, 1(??):266, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900321.

Kesteven:1981:BRF

[Kes82c] G. L. Kesteven. Book review: Food from the sea: James Nicolson. Cassell, London, 1979. 205 pp., £6.50, ISBN 0-304-30315-1. Fisheries Research, 1(??):72–73, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900096.

Kesteven:1981:BRS

[Kes82d] G. L. Kesteven. Book review: Study of the sea: E. M. Thomasson (Editor). Fishing News Books, Farnham, Surrey, Great Britain, 1981. 256 pp., 68 illus., £25.00, ISBN 0-85238-112-3. Fisheries Research, 1(??):261–262, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900291.

Kesteven:1981:IIF

[Kes82e] G. L. Kesteven. The international issue of fisheries research. Fisheries Research, 1(??):191–194, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900242.

Kesteven:1981:SRF

[Kes82f] G. L. Kesteven. Scientific research for the fisheries of today and tomorrow. Fisheries Research, 1(??):1–2, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900023.

Kesteven:1983:BRF

[Kes83a] G. L. Kesteven. Book review: Fish by-catch… Bonus from the sea: FAO/IDRC. IDRC-198e, International Development Research Centre, Ottawa, Canada, 1982, ISBN 0-88936-336-6. Fisheries Research, 2(1):61–63, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901030.

Kesteven:1983:BRP

[Kes83b] G. L. Kesteven. Book review: Principles of fishery science: W. Harry Everhart and William D. Youngs. Cornell University Press, Ithaca and London, 2nd edn. 1981. 349 pp., 67 illus., US$17.50./£ 22.70, ISBN 0-8014-1334-6. Fisheries Research, 2(2): 159, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683901169.

Kesteven:1983:BRW

[Kes83c] G. L. Kesteven. Book review: Words of the Lagoon: Fishing and marine lore in the Palau district of Micronesia: R. E. Johannes. University of California Press, Berkeley, Los Angeles, London, 1981, 245 pp., ISBN 0-520-03929-7. Fisheries Research, 2(1): 70–72, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683901066.

Kesteven:1984:BRF

[Kes84a] G. L. Kesteven. Book review: Fishery Economics and Management in India: P. S. Rao. Pioneer Publishers and Distributors, Bombay, 1983, 119 tables, 27 plates, 10 figs., 356 pp., Rs. 225.00, US $25.00. Fisheries Research, 2(4):318–319, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900353.

Kesteven:1984:BRMa

[Kes84b] G. L. Kesteven. Book review: Managing the ocean resources of the United States: the role of the federal marine sanctuaries program: By D. P. Finn. Springer-Verlag, Berlin, Heidelberg, New York, 1982, ix + 192 pp., paperback price DM 38.00 (approx. US$15.90), ISBN 3-540-11583-8. Fisheries Research, 2(3): 234–236, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900122.

Kesteven:1984:BRMb

[Kes84c] G. L. Kesteven. Book review: Memoria, 1er coloquio international sobre legislacion pesquero: Volumen 5. Departmento de Pesca, Institute de Investigaciones Juridicas de la UNAM, Mexico, 1981, 164 pp., US$7.50. Fisheries Research, 2(3):236– 237, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900134.

Kesteven:1984:BRP

[Kes84d] G. L. Kesteven. Book review: Pacific coast clam fisheries: T. D. Schink, K. A. McGraw and K. Chew. Washington Sea Grant Communications, University of Washington, 1983, 16 tables, 19 figs., iv + 72 pp., price $4.50 (US funds). Fisheries Research, 2 (4):319–320, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900365.

Kesteven:1985:BRM

[Kes85a] G. L. Kesteven. Book review: Marine resource economics: Volume 1, Number 1. Jon G. Sutinen (Editor). Crane Russak, New York, 1984, 115 pp., US $78.00, ISSN 0738-1360. Fisheries Research, 3(??):174–177, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783685900189.

Kesteven:1985:BRS

[Kes85b] G. L. Kesteven. Book review: Scallop and queen fisheries in the British Isles: by James Mason. (A Buckland Foundation Book) Fishing News Books, Farnham, Surrey, Gt. Britain, 144 pp., 37 figs., 23 plates, 8 tables, £6.90, ISBN 0-85238-128-X. Fisheries Research, 3(??):178, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783685900190.

King:1987:APG

[KFM87] D. P. F. King, A. Ferguson, and I. J. J. Moffett. Aspects of the population genetics of herring, Clupea harengus, around the British Isles and in the Baltic Sea. Fisheries Research, 6(1):35– 52, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900051.

Kirkegaard:1983:BRP

[Kir83] Ian Kirkegaard. Book review: Proceedings of the International Shrimp Releasing, Marking and Recruitment Workshop, 25th–29th November 1978, Salmiya, State of Kuwait: Edited by A. S. D. Farmer. Kuwait Institute for Scientific Research, October 1981, 415pp. Fisheries Research, 2(1):78–80, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783683901108.

Kirkwood:1986:BRR

[Kir86] G. P. Kirkwood. Book review: Resource management proceedings, Ashland, Oregon, 1984: M. Mangel (Editor) Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1985, 138 pp., price DM25.00, ISBN 3-540-15982-7. Fisheries Research, 4(3–4):362– 363, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900196.

Lander:1981:LTB

[Lan82] Robert H. Lander. A life table and biomass estimate for Alaskan fur seals. Fisheries Research, 1(??):55–70, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900072.

Lanzing:1985:BRH

[Lan85] W. J. R. Lanzing. Book review: Histology of the striped bass: by David B. Groman. American Fisheries Society, Bethesda, MD, 1982, Monograph No. 3, 116 pp., price $11.00, ISBN 0-913235-14-8, ISSN 0097-0638. Fisheries Research, 3(??):376– 377, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900372.

Lee:1988:BRC

[Lee88] David J. Lee. Book review: Carbon dynamics in eutrophic, temperate lakes: by Bo Riemann and Morten Sondergaard (Editors). Elsevier Science Publishing Company Inc., Amsterdam, 1986, 284 pages, $66.75./Dfl. 150.00, ISBN:0-444-42736-8. Fisheries Research, 6(4):396–397, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783688900094.

Lewis:1985:BRP

[Lew85] Marlon R. Lewis. Book review: Polarographic oxygen sensors: Aquatic and physiological applications: E. Gnaiger and H. Forstner (Editors). Springer-Verlag, Berlin, Heidelberg, New York, 1983, 142 figs., xii + 370 pp., cloth DM 118.00, approx. US $50.90, ISBN 3-540-11654-0. Fisheries Research, 3(??):173– 174, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900177.

Lai:1987:EAE

[LG87] Han Lin Lai and Donald R. Gunderson. Effects of ageing errors on estimates of growth, mortality and yield per recruit for walleye pollock (Theragra chalcogramma). Fisheries Research, 5(2–3): 287–302, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900488.

Lindqvist:1987:BRF

[Lin87] O. V. Lindqvist. Book review: Fish physiology: Recent advances: Stefan Nilsson and Susanne Holmgren (Editors). Croom Helm Publishers, Beckenham, 1986, 198 pp., ill., £27.50, ISBN 0-7099-1837-2. Fisheries Research, 5(4):418–419, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900592.

Liston:1981:BRT

[Lis82] John Liston. Book review: Trends in fish utilization: J. J. Connell and R. Hardy. Fishing News Books Ltd., Farnham, Surrey, Great Britain, 1982. 101 pp., 44 illus., 216 × 138 mm, £6.00 + post/packing, ISBN 0-85238-120-4. Fisheries Research, 1(??):

345–346, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900400.

Lockwood:1988:BRD

[Loc88] Stephen J. Lockwood. Book review: Developments in fisheries research in Scotland: by R. S. Bailey and B. B. Parrish (Editors). Fishing News Books Ltd., Farnham, Surrey, 1987, 282 pp., ill., £27.50, ISBN: 0-85238-144-1. Fisheries Research, 6(3): 301–303, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900227.

Longhurst:1985:BRN

[Lon85] A. R. Longhurst. Book review: North sea dynamics: J. Sündermann and W. Lenz (Editors). Springer-Verlag, Berlin, Heidelberg, New York, 1983, 261 figs., 693 pp., cloth DM 98.00, approx. US $42.30, ISBN 3-540-12013-0. Fisheries Research, 3(??): 82–84, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900104.

Lockwood:1983:ECM

[LPE83] S. J. Lockwood, M. G. Pawson, and D. R. Eaton. The effects of crowding on mackerel (Scomber scombrus L.) — physical condition and mortality. Fisheries Research, 2(2):129–147, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683901145.

Lagardere:1989:CSA

[LS89] J. P. Lagardère and D. Sureau. Changes in the swimming activity of the sole (Solea vulgaris Quensel, 1806) in relation to winter temperatures in a saltmarsh: Observations using ultrasonic telemetry. Fisheries Research, 7(3):233–239, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578368990057X.

Liu:1985:TCE

[LSC85] Hsi-Chiang Liu, K. J. Sainsbury, and Tai-Sheng Chiu. Trawl codend mesh selectivity for some fishes of North-Western Australia. Fisheries Research, 3(??):105–129, ???? 1985. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578368590013X.

Lleonart:1985:CEV

[LSM85] Jordi Lleonart, Jordi Salat, and Enrique Macpherson. CVPA, an expanded VPA with cannibalism. Application to a hake population. Fisheries Research, 3(??):61–79, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900086.

Leslie:1989:DNM

[LT89] J. K. Leslie and C. A. Timmins. Double nets for mesh aperture selection and sampling in ichthyoplankton studies. Fisheries Research, 7(3):225–232, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900568.

MacLennan:1981:DFP

[Mac82] D. N. MacLennan. The drag of four-panel demersal trawls. Fisheries Research, 1(??):23–33, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900047.

MacLennan:1983:BRR

[Mac83] D. N. MacLennan. Book review: Recent developments in side scan sonar techniques: W. G. A. Russell-Cargill (Editor). Published by the editor, University of Capetown, South Africa, 1982, 141 pp., US$16.00, ISBN: 0-7992-0465-X. Fisheries Research, 2 (2):160–162, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901170.

Magnussen:1986:BRS

[Mag86] Ola M. Magnussen. Book review: Shipboard refrigeration and fish processing: by N. G. Kondrashova. A. A. Balkema, Rotterdam, 1984, 250 pp., $24.00, £15.00, Dfl.68.25, ISBN 90-6191-432-9. Fisheries Research, 4(3–4):359–360, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900172.

Mukundan:1986:RAF

[MAN86] M. K. Mukundan, P. D. Antony, and M. R. Nair. A review on autolysis in fish. Fisheries Research, 4(3–4):259–269, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578368690007X.

Marais:1985:SFI

[Mar85] J. F. K. Marais. Some factors influencing the size of fishes caught in gillnets in eastern Capeo estuaries. Fisheries Research, 3(??): 251–261, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900268.

Matuda:1988:HHB

[Mat88] K. Matuda. Headline height of bottom gill nets set across a water flow. Fisheries Research, 6(2):167–179, January 1988. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783688900355.

McIntyre:1987:SRF

[McI87] A. D. McIntyre. The scope of research on fisheries. Fisheries Research, 6(1):1–3, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900026.

McIntyre:1989:BRA

[McI89a] Alasdair D. McIntyre. Book review: Asian fisheries science: Vol. 1, No. 1, December 1987. Published by the Asian Fisheries Society, Manila, The Philippines, 106 pp., annual subscription for one volume (two issues) US$9.00, ISSN: 0116-6514. Fisheries Research, 7(4):385–386, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900702.

McIntyre:1989:BRS

[McI89b] Alasdair D. McIntyre. Book review: Smiths’ sea fishes: Margaret M. Smith and Phillip C. Heemstra (Editors), Springer-Verlag, 1986, 1047 pp., DM 198, ISBN 3-540-16851-6. Fisheries Research, 7 (4):385, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900696.

Manooch:1987:AGY

[MD87] Charles S. Manooch and Caroline L. Drennon. Age and growth of yellowtail snapper and queen triggerfish collected from the U.S. Virgin Islands and Puerto Rico. Fisheries Research, 6(1):53– 68, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900063.

Mitchell:1981:BRI

[Mit82a] C. L. Mitchell. Book review: International joint ventures in world fisheries: V. Kaczynski and D. LeVieil. Technical Report, Division of Marine Resources, University of Washington. Sea Grant, Washington, DC, 1980, 33 pp., US $2.50. Fisheries Research, 1 (??):74–75, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900102.

Mitchell:1981:BRF

[Mit82b] Carlyle L. Mitchell. Book review: Food from the sea: the economics and politics of ocean fisheries: F. W. Bell. Westview Student Edition, Westview Press, Boulder, CO, 1978. Distributor for Europe: Bowker Publishing Co., Essex, Great Britain. 380 pp., US $17.50 paperback/US $43.00 hardback, ISBN 0-89158-353-X. Fisheries Research, 1(??):348–349, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900424.

Mitchell:1984:BRE

[Mit84] C. L. Mitchell. Book review: Economies of Ocean Resources: G. M. Brown, Jr. and J. A. Crutchfield, Editors. A Washington Sea Grant Publication, University of Washington Press, Seattle, WA, U.S.A., 1982, pp. 242, US$12.00, ISBN 0-295-95982-7. Fisheries Research, 2(3):233–234, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900110.

Mitchell:1985:BRL

[Mit85] C. L. Mitchell. Book review: The legal regime of fisheries in the Caribbean region: by W. R. Edeson and J. F. Pulvenis. Lecture Notes on Coastal and Estuarine Studies, Vol. 7. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1983, x + 204 pp., soft cover DM 43.00, approx. US $16.70, ISBN 3-540-12698-8. Fisheries Research, 3(??):380–381, ???? 1985. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900414.

Mitchell:1987:BRE

[Mit87] C. L. Mitchell. Book review: The economics of fisheries management: Lee G. Anderson. The John Hopkins University Press, Baltimore, MD 1986, vii–xx 296 pp., $32.95, ISBN 0-8018-3253-5. Fisheries Research, 5(4):414–416, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900579.

Mitchell:1988:BRS

[Mit88] C. L. Mitchell. Book review: Salmon: Economics and marketing: by Susan A. Shaw and James F. Muir. Croom Helm, London; Timber Press, Portland, OR, 1987, 1987, 270 pp., ill., £25.00, ISBN:0-7099-3344-4. Fisheries Research, 6(4):393–394, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900070.

Marshall:1988:PRB

[ML88] B. E. Marshall and J. D. Langerman. A preliminary reappraisal of the biomass of inshore fish stocks in Lake Kariba. Fisheries Research, 6(2):191–199, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900379.

Mason:1985:AGM

[MM85] Diane L. Mason and Charles S. Manooch. Age and growth of mutton snapper along the east coast of Florida. Fisheries Research, 3(??):93–104, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783685900128.

Mabawonku:1986:RPI

[MO86] A. F. Mabawonku and S. A. Olomola. Regulations and policies on the importation of fishing materials in Nigeria: 1970–1980. Fisheries Research, 4(3–4):213–222, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900032.

Mohr:1983:BRI

[Moh83] Viggo Mohr. Book review: Introduction to fishery by-products: Malcolm Windsor and Stuart Barlow, 187 pp., Fishing News Books, Surrey, U.K., £13.50 + £1.35 (p+p) ISBN 0-85238-115-8. Fisheries Research, 2(1):72–74, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901078.

Moriarty:1986:RMY

[Mor86] Christopher Moriarty. Riverine migration of young eels Anguilla anguilla (L.). Fisheries Research, 4(1):43–58, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900275.

Moses:1988:GMP

[Mos88] Boniface S. Moses. Growth, mortality and potential yield of bonga, Ethmalosa fimbriata (Bowdich 1825) of Nigerian inshore waters. Fisheries Research, 6(3):233–247, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900161.

Meenakumari:1985:SMT

[MR85] B. Meenakumari and K. V. Mohan Rajan. Studies on materials for traps for spiny lobsters. Fisheries Research, 3(??):309– 321, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900293.

Main:1985:TET

[MS85] J. Main and G. I. Sangster. Trawling experiments with a two-level net to minimise the undersized gaboid by-catch in a Nephrops fishery. Fisheries Research, 3(??):131–145, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900141.

Macer:1987:ESF

[MS87] C. T. Macer and J. G. Shepherd. The effect of species and fishery interactions on the assessment of some North Sea gadoids. Fisheries Research, 5(2–3):303–317, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578368790049X.

McShane:1989:DMF

[MS89] Paul E. McShane and Murray G. Smith. Direct measurement of fishing mortality in abalone (Haliotis rubra Leach) off southeastern Australia. Fisheries Research, 8(2):93–102, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900246.

Miller:1983:DAG

[MWS83] David R. Miller, John G. Williams, and Carl W. Sims. Distribution, abundance and growth of juvenile salmonids off the coast of Oregon and Washington, summer (1980). Fisheries Research, 2 (1):1–17, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683900991.

Moring:1986:MPC

[MYH86] John R. Moring, Ronald L. Youker, and Robert M. Hooton. Movements of potamodromous coastal cutthroat trout, Salmo clarki clarki, inferred from tagging and scale analyses. Fisheries Research, 4(3–4):343–354, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900135.

Naevdal:1986:BRO

[Nae86] G. Naevdal. Book review: Oberservations on the ecology of Western Cape Cod Bay, Massachusetts: J. D. Davis and D. Merriman (Editors). Springer-Verlag, Heidelberger Platz 3, Postfach D-1000, Berlin 33, 1984, 289 pp., price DM58.00, ISBN 3-540-96084-8. Fisheries Research, 4(2):174–176, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900469.

Nakken:1985:BRF

[Nak85] Odd Nakken. Book review: Fisheries sonar: by R. B. Mitson. Fishing News Books, Surrey, England, 1983, 288 pp., price £26.00, ISBN 0-85238-124-7. Fisheries Research, 3(??):81–82, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900098.

Nauen:1985:BRA

[Nau85] Cornelia E. Nauen. Book review: Acid rain/fisheries: Proceedings of an International Symposium on Acidic Precipitation and Fishery Impacts in Northeastern North America, Cornell University, Ithaca, NY, 2–5 August 1981. Compiled under the direction of Terry A. Haines, Chairman, and Raymond E. Johnson (Editor). Published by the Northeastern Division, American Fisheries Society, Bethesda, MD 20814, 1982, 357 pp., illustrated, U.S. $28.00, LCCCN 82-74271. Fisheries Research, 3(??):384– 386, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900438.

Newland:1989:SOB

[NC89] P. L. Newland and C. J. Chapman. The swimming and orientation behaviour of the Norway lobster, Nephrops norvegicus (L.), in relation to trawling. Fisheries Research, 8(1):63– 80, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900416.

Nasir:1989:DLW

[NNAS89] N. A. Nasir, A. K. Naama, and A. Al-Saboonchi. The distribution, length–weight relationships, food and feeding of the cyprinid fish Barbus sharpeyi from Al-Hammar Marsh, Iraq. Fisheries Research, 7(1–2):175–181, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900143.

Nakken:1987:DGN

[NR87] Odd Nakken and Askjell Raknes. The distribution and growth of Northeast Arctic cod in relation to bottom temperatures in the Barents Sea, 1978–1984. Fisheries Research, 5(2–3):243– 252, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900440.

Orlowski:1989:SFB

[Orł89] A. Orłowski. Seasonal fluctuations of biomass distribution based on results of hydroacoustic surveys of the Polish fishery zone. Fisheries Research, 8(1):25–34, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900386.

Pak:1989:MMD

[Pak89] Sergey B. Pak. Mathematical model of dynamics of spawning sizestructured population (with an application). Fisheries Research, 8 (2):141–158, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900283.

Papaconstantinou:1984:AGY

[Pap84] C. Papaconstantinou. Age and growth of the yellow gurnard (Trigla lucerna L. 1758) from the Thermaikos Gulf (Greece) with some comments on its biology. Fisheries Research, 2(4):243–255, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783684900286.

Pawson:1989:BRI

[Paw89] M. G. Pawson. Book review: Indonesian marine capture fisheries: C. Bailey, A. Dwiponggo and F. Marahudin, International Center for Living Aquatic Resources Management (ICLARM), Manila, Philippines, available from ICLARM, MC P. O. Box 1501, Makati, Metro Manila, Philippines as well as from TRIPS (Tropical Scientific) Books, Raiffenstrasse 24, D-6070 Langen, F.R.G., xi + 196 pp. US $9.50 (surface), US $18 (airmail), ISBN: 971-10-2230-3. Fisheries Research, 7(1–2):185–186, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900179.

Paulovits:1986:ADG

[PB86] G. Paulovits and P. Biró. Age determination and growth of eel, Anguilla anguilla (L.), in Lake Fertő, Hungary. Fisheries Research, 4(2):101–110, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900366.

Papaconstantinou:1989:FIB

[PC89] C. Papaconstantinou and E. Caragitsou. Feeding interaction between two sympatric species Pagrus pagrus and Phycis phycis around Kastellorizo Island (Dodecanese, Greece). Fisheries Research, 7(4):329–342, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900659.

Patrick:1985:RFS

[PCS^{+}85] Paul H. Patrick, A. E. Christie, D. Sager, C. Hocutt, and J. Stauffer. Responses of fish to a strobe light/air-bubble barrier. Fisheries Research, 3(??):157–172, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900165.

Pringle:1986:CET

[PDS86] J. D. Pringle, R. E. Duggan, and G. J. Sharp. A cost evaluation of techniques designed to assess Canadian lobster fishing effort. Fisheries Research, 4(3–4):283–295, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900093.

Peeling:1984:BRS

[Pee84] D. A. Peeling. Book review: Squid Jigging from Small Boats: By Mototsuqu Hamabe, Chikamasa Hamuro and Michio Ogura. Fishing News Books, Farnham, Surrey, U. K., 1982, A F.A.O. Fishing Manual, 84 pp., price £6.00 plus 60p post/packing, 42 illus., ISBN 0-85238-122-0. Fisheries Research, 2(3):227, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900080.

Petr:1986:BRF

[Pet86] T. Petr. Book review: Freshwater fisheries management: by Robin G. Templeton (Editor). Fishing News Books, 1 Long Garden Walk, Farnham, Surrey, England, 1984, 192 pp., Softcover £12.50, ISBN 0 85238 130 1. Fisheries Research, 4(3–4):361– 362, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900184.

Pillay:1986:BRP

[Pil86] T. V. R. Pillay. Book review: Pond fisheries: by F. G. Martyshev. A. A. Balkema, Rotterdam, 1983, 454 pp., price Dfl. 75.00, U.S. $28.00, £17.50. Fisheries Research, 4(1):80– 81, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900329.

Penczak:1985:FPM

[PJDZ85] T. Penczak, J. Jankov, T. S. J. Dikov, and M. Zalewski. Fish production in the Mesta River, Rila Mountain, Samokov, Bulgaria. Fisheries Research, 3(??):201–221, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900220.

Phillips:1986:TSA

[PM86] A. C. Phillips and J. C. Mason. A towed, self-adjusting sled sampler for demersal fish eggs and larvae. Fisheries Research, 4(3– 4):235–242, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900056.

Pollard:1983:BRB

[Pol83] D. A. Pollard. Book review: Biology of Fishes: Q. Bone and N. B. Marshall. Tertiary Level Biology Series. Blackie and Son, Glasgow and London (distributed in U.S.A. by Chapman and Hall, New York), 1982. ix + 253 pp., paperback, price: £11.95. ISBN: 0-216-91017-X. Fisheries Research, 2(2):163–166, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783683901194.

Priede:1984:BSC

[Pri84a] I. G. Priede. A basking shark (Cetorhinus maximus) tracked by satellite together with simultaneous remote sensing. Fisheries Research, 2(3):201–216, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900031.

Priede:1984:BRC

[Pri84b] I. G. Priede. Book review: Chemoreception in fishes: T. J. Hara (Editor). Developments in Aquaculture and Fisheries Science No. 8, Elsevier Science Publishers, Amsterdam, New York, 1982, x + 434 pp., price US$91.50 Dfl. 215.00, ISBN 0-444-42040-1. Fisheries Research, 2(3):222–223, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900067.

Priede:1986:BRC

[Pri86] I. G. Priede. Book review: The control of fish migration: Reginald J. F. Smith, Zoophysiology. Vol. 17. Springer-Verlag, Berlin, 1985, 243 pp., DM 136.00, ISBN 3-540-13707-6. Fisheries Research, 4(2):169–171, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900433.

Priede:1989:BRS

[Pri89] I. G. Priede. Book review: The sei whale: Population biology, ecology and management: J. W. Horwood. Croom Helm, Beckenham, Kent. 1987, 375 pp., £29.50, ISBN: 0-7099-4786-0. Fisheries Research, 7(1–2):183–184, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900155.

Patterson:1985:OBL

[PW85] R. N. Patterson and K. C. Watts. The otter board as a lowaspect-ratio wing at high angles of attack; some theoretical aspects. Fisheries Research, 3(??):351–372, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900335.

Patterson:1986:OBL

[PW86] R. N. Patterson and K. C. Watts. The otter board as a lowaspect-ratio wing at high angles of attack; an experimental study. Fisheries Research, 4(2):111–130, July 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900378.

Quinn:1985:LTE

[Qui85] Terrance J. Quinn. Line transect estimators for schooling populations. Fisheries Research, 3(??):183–199, ???? 1985. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783685900219.

Rossi:1984:EFM

[RC84] Remigio Rossi and Alessandra Cannas. Eel fishing management in a hypersaline lagoon of southern Sardinia. Fisheries Research, 2 (4):285–298, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900316.

Ranta:1989:PLS

[RL89] Esa Ranta and Kai Lindström. Prediction of lake-specific fish yield. Fisheries Research, 8(2):113–128, November 1989. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
016578368990026X.

Rudstam:1988:DST

[RLH88] L. G. Rudstam, T. Lindem, and S. Hansson. Density and in situ target strength of herring and sprat: a comparison between two methods of analyzing single-beam sonar data. Fisheries Research, 6(4):305–315, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368890001X.

Rosenberg:1981:CHS

[RP82] Rutger Rosenberg and Lars-Erik Palmén. Composition of herring stocks in the Skagerrak–Kattegat and the relations of these stocks with those of the North Sea and adjacent waters. Fisheries Research, 1(??):83–104, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900126.

Ruckes:1986:BRM

[Ruc86] E. Ruckes. Book review: Marketing in fisheries and aquaculture: by Ian Chaston. Fishing News Books, Farnham, Surrey, Gt. Britain, 1983, 144 pp., 31 figs., 4 tables, price £6.50, ISBN 0-85238-129-8. Fisheries Research, 4(1):75–76, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900299.

Saetersdal:1981:BRP

[Sae82] Gunnar Saetersdal. Book review: The politics of Pacific island fisheries: George Kent. Westview Replica Edition, Westview Press, Boulder, CO, 1980. Distributor for Europe: Bowker Publishing Co., Essex, Great Britain, 191 pp., £14.25, ISBN 0-89158-683-0. Fisheries Research, 1(??):350–352, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900436.

Sanders:1988:MPN

[San88] M. J. Sanders. Mean population number and the De Lury and Leslie methods. Fisheries Research, 6(2):153–165, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783688900343.

Smith:1989:CEC

[SBPR89] M. P. Lincoln Smith, J. D. Bell, D. A. Pollard, and B. C. Russell. Catch and effort of competition spearfishermen in Southeastern Australia. Fisheries Research, 8(1):45–61, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900404.

Schumacher:1987:DPB

[Sch87] J. D. Schumacher. Description of the physical and biological environment Rapporteur’s report. Fisheries Research, 5(2–3):111– 118, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900361.

Sirisena:1989:NFR

[SD89] H. K. G. Sirisena and Sena S. De Silva. New fish resources of reservoirs in Sri Lanka. II. Further studies on a gillnet fishery for minor cyprinids. Fisheries Research, 7(1–2):17–29, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900040.

Stewart:1985:MGN

[SF85] P. A. M. Stewart and R. S. T. Ferro. Measurements on gill nets in a flume tank. Fisheries Research, 3(??):29–46, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900062.

Stewart:1987:FEI

[SF87] P. A. M. Stewart and R. S. T. Ferro. Four experiments investigating codend drag. Fisheries Research, 5(4):349–358, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368790052X.

Sause:1987:LSJ

[SGB87] B. L. Sause, D. Gwyther, and D. Burgess. Larval settlement, juvenile growth and the potential use of spatfall indices to predict recruitment of the scallop Pecten alba Tate in Port Phillip Bay, Victoria, Australia. Fisheries Research, 6(1):81– 92, November 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900087.

Shelton:1981:CES

[SH82] R. G. J. Shelton and W. B. Hall. A comparison of the efficiency of the Scottish creel and the inkwell pot in the capture of crabs and lobsters. Fisheries Research, 1(??):45–53, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900060.

Sager:1987:EFR

[SHS87] David R. Sager, Charles H. Hocutt, and Jay R. Stauffer. Estuarine fish responses to strobe light, bubble curtains and strobe light/bubble-curtain combinations as influenced by water flow rate and flash frequencies. Fisheries Research, 5(4):383–399, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900543.

Smith:1986:SDH

[SJ86] P. J. Smith and A. Jamieson. Stock discreteness in herrings: a conceptual revolution. Fisheries Research, 4(3–4):223–234, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900044.

Surendran:1989:SSC

[SJS^{+}89] P. K. Surendran, Jose Joseph, A. V. Shenoy, P. A. Perigreen, K. Mahadeva Iyer, and K. Gopakumar. Studies of spoilage of commercially important tropical fishes under iced storage. Fisheries Research, 7(1–2):1–9, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900027.

Saetersdal:1987:EAR

[SL87] Gunnar Sætersdal and Harald Loeng. Ecological adaptation of reproduction in Northeast Arctic cod. Fisheries Research, 5(2–3): 253–270, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900452.

Sloan:1987:RCC

[SLH87] N. A. Sloan, C. P. Lauridsen, and R. M. Harbo. Recruitment characteristics of the commercially harvested red sea urchin Strongylocentrotus franciscanus in Southern British Columbia, Canada. Fisheries Research, 5(1):55–69, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783687900154.

Smith:1981:BRF

[Smi82] Paul E. Smith. Book review: Fisheries oceanography and ecology: Taivo Laevastu and Murray L. Hayes. Fishing News Books Ltd., Farnham, Surrey, Great Britain. 199 pp., 81 illus., 246 × 189 mm, £19.50 + £1.95 post/packing, ISBN 0-85238-117-4. Fisheries Research, 1(??):346–348, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900412.

Smith:1983:BRH

[Smi83] Lynwood S. Smith. Book review: Hearing and sound communication in fishes: Edited by W. N. Tavolga, A. N. Popper and R. R. Fay. Springer-Verlag, Berlin and New York, 1981, 196 figs, XVI + 608 pp., DM 118, approx. US $55.00, ISBN 3-540-90590-1. Fisheries Research, 2(1):77–78, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783683901091.

Sandlund:1989:IPG

[SN89] Odd Terje Sandlund and Tor F. Næsje. Impact of a pelagic gillnet fishery on the polymorphic whitefish (Coregonus lavaretus L. s.l.) population in Lake Femund, Norway. Fisheries Research, 7 (1–2):85–97, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368990009X.

Somers:1986:BRE

[Som86] I. F. Somers. Book review: Exploitation of marine communities: R. M. May (Editor). Dahlem Konferenzen 1984, Springer-Verlag, Berlin, New York, Tokyo, 370 pp., DM 48.00, ISBN 3-540-15028-5. Fisheries Research, 4(3–4):356–358, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900159.

Somers:1987:BRS

[Som87] I. F. Somers. Book review: Systems analysis and simulation in wildlife and fisheries sciences: William E. Grant. Wiley, Chichester, 1986, 338 pp. ill., £45.65, ISBN 0-471-89236-X. Fisheries Research, 5(4):413–414, August 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783687900567.

Somers:1988:BRF

[Som88] Ian Somers. Book review: Fisheries development: 2000 A.D., by K. K. Trivedi (Editor). A. A. Balkema Publishers, Rotterdam, The Netherlands, 1986, 274 pp., figs., tables, cloth, Dfl. 85/$34/£26.50, ISBN:90-6191-484-1. Fisheries Research, 6(4):394– 395, September 1988. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783688900082.

Somers:1989:BRP

[Som89] I. F. Somers. Book review: The Peruvian anchoveta and its upwelling ecosystem: D. Pauly and I. Tsukayama (Editors), ICLARM Studies and Reviews 15, Instituto del Mar del Peru (IMARPE), Deutsche Gesellschaft Für Technische Zusammenarbeit (GTZ) and International Center for Living Aquatic Resources Management (ICLARM). Available from ICLARM, MC P. O. Box 1501, Makati, Metro Manilla, Philippines, TRIOPS (Tropical Scientific Books), Raiffenstrasse 24, D-6070 Langen, F.R.G., xii + 351 pp., hardcover price US $17.50 (surface) $36.00 (airmail), ISBN: 971-10-2234-6. Fisheries Research, 7(1–2):187–188, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900180.

Summers:1981:SMW

[SRRU82] J. Kevin Summers, William A. Richkus, Richard N. Ross, and Robert E. Ulanowicz. A statistical model of white perch (Morone americana) harvest assessment in relation to environmental variation. Fisheries Research, 1(??):289–298, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900369.

Sumpton:1989:ICM

[SSR89] W. D. Sumpton, P. J. Smith, and B. G. Robotham. The influence on catch of monofilament and multifilament netting in otter prawn trawls. Fisheries Research, 8(1):35–44, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900398.

Stober:1981:RCI

[ST82] Q. J. Stober and R. W. Tyler. Rule curves for irrigation drawdown and Kokanee salmon (Oncorhynchus nerka) egg to fry survival. Fisheries Research, 1(??):195–218, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783681900254.

Suboski:1989:LST

[ST89] Milton D. Suboski and Jennifer J. Templeton. Life skills training for hatchery fish: Social learning and survival. Fisheries Research, 7(4):343–352, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900660.

Stewart:1981:IRF

[Ste82] Peter A. M. Stewart. An investigation into the reactions of fish to electrified barriers and bubble curtains. Fisheries Research, 1 (??):3–22, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900035.

Stewart:1988:MET

[Ste88] P. A. M. Stewart. Measurements of the effects of tidal flow on the headline heights of bottom-set gillnets. Fisheries Research, 6(2):181–189, January 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783688900367.

Suomala:1981:BRE

[Suo82] J. B. Suomala. Book review: Echo sounding and sonar for fishing: (Food and Agricultural Organization Fishing Manuals) Fishing News Books Ltd., Surrey, Great Britain. 104 pp., 120 illus., 216 × 138 mm, £6.50, ISBN 0-85238-110-7. Fisheries Research, 1 (??):177–178, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900199.

Skurdal:1985:CSO

[SVQ85] Jostein Skurdal, Leif Asbjørn Vøllestad, and Tore Qvenild. Comparison of scales and otoliths for age determination of whitefish Coregnous lavaretus. Fisheries Research, 3(??):237–243, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783685900244.

Siaw-Yang:1984:CAF

[SY84] Yap Siaw-Yang. Cohort analysis on a freshwater fish Osteochilus hasselti C. & V. (Cyprinidae) at Bukit Merah Reservoir, Malaysia. Fisheries Research, 2(4):299–314, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783684900328.

Tang:1985:MRS

[Tan85] Qisheng Tang. Modification of the Ricker stock recruitment model to account for environmentally induced variation in recruitment with particular reference to the blue crab fishery in Chesapeake Bay. Fisheries Research, 3(??):13–21, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783685900049.

Tweddle:1988:CEB

[TB88] D. Tweddle and P. Bodington. A comparison of the effectiveness of black and white gillnets in Lake Malawi, Africa. Fisheries Research, 6(3):257–269, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783688900185.

Tomasson:1985:DML

[TBH85] T. Tomasson, M. N. Bruton, and K. C. D. Hamman. The demography and management of large cyprinids in a reservoir on the Orange River, South Africa. Fisheries Research, 3(??):279– 308, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900281.

Thorogood:1986:ARB

[Tho86a] John Thorogood. Aspects of the reproductive biology of the southern bluefin tuna (Thunnus maccoyii). Fisheries Research, 4(3– 4):297–315, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016578368690010X.

Thorogood:1986:BRF

[Tho86b] John Thorogood. Book review: Fisheries techniques: Larry A. Nielsen and David L. Johnson (Editors). American Fisheries Society, 5410 Grosvenor Lane, Bethesda, MD 20814, U.S.A., 1983, 468 pp., price US $32.00, ISBN 0-913235-00-8. Fisheries Research, 4(1):84–85, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783686900342.

Thorsteinsson:1988:BRC

[Tho88] Gudni Thorsteinsson. Book review: Calculations for fishing gear designs: by A. L. Fridman, revised, edited and enlarged by P. J. G. Carrothers. An F.A.O. Fishing Manual. Fishing News Books Ltd, Farnham, Surrey, 1986, 242 pp, 104 figs. ISBN 0-85238-141-7. Fisheries Research, 6(3):299–301, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783688900215.

Thorsteinsson:1989:BRO

[Tho89a] Gudni Thorsteinsson. Book review: Ownership and productivity of marine fishery resources: Elmer A. Keen, McDonald and Woodward Publishing Co., Blacksburg, VA, 1988, 122 pp., US$10.95 (for the U.S. direct orders should be prepared, including US$1.80 shipping charge for first copy and US$0.30 for each additional copy; countries other than the U.S. should include US$1.50 for international shipping charges), ISBN 0-939923-05-X. Fisheries Research, 8(2):198–200, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900325.

Thorsteinsson:1989:BRL

[Tho89b] Gudni Thorsteinsson. Book review: The lobster gangs of Maine: James M. Acheson, University Press of New England, Hanover and London, 1988, 181 pp., £6.50, ISBN 0-87451-451-7. Fisheries Research, 8(2):197–198, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900313.

Thresher:1984:BRL

[Thr84] R. E. Thresher. Book review: The Larvae of Indo–Pacific Coral Reef Fishes: by J. M. Leis and D. S. Rennis. University N.S.W. Press, Sydney, and University Hawaii Press, Hawaii, 1983, 269 pp., 75 figs., $A 23.95, ISBN 0-86840-129-3. Fisheries Research, 2(4):320–321, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900377.

Thresher:1985:BRC

[Thr85] Ronald E. Thresher. Book review: Carribean coral reef fishery resources: J. L. Munro (Editor). ICLARM Studies and Reviews No. 7, ICLARM, Metro Manila, Philippines, 1983, 276 pp., price US $16.00 surface, US $33.00 airmail (paper), US $19.50 surface, US $37.00 airmail (cloth), ISBN 971-10-2200-1 cloth, ISBN 971-10-2201-X paperback. Fisheries Research, 3(??):373– 374, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900347.

Tom:1983:ABL

[TL83] Moshe Tom and Chanan Lewinsohn. Aspects of the benthic life cycle of Penaeus (Melicertus) japonicus Bate (Crustacea Decapoda) along the south-eastern coast of the Mediterranean. Fisheries Research, 2(2):89–101, September 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783683901121.

Tinsley:1989:PSS

[TNW89] V. Randall Tinsley, Larry A. Nielsen, and David H. Wahl. Pushnet sampling as a supplement to seine sampling in rivers. Fisheries Research, 7(3):201–206, May 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900544.

Troadec:1981:BRR

[Tro82] J.-P. Troadec. Book review: A research framework for traditional fisheries: Ian R. Smith. ICLARM, Metro Manila, The Philippines, 1979. ICLARM Studies and Reviews, No. 2. ix + 40 pp., paperback, ISSN 0115-4389. Fisheries Research, 1(??):185–187, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900229.

Tsimenides:1989:ADG

[TT89] N. Tsimenides and G. Tserpes. Age determination and growth of swordfish Xiphias gladius L., 1758 in the Aegean Sea. Fisheries Research, 8(2):159–168, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900295.

Torricelli:1981:MGM

[TTA82] P. Torricelli, P. Tongiorgi, and P. Almansi. Migration of grey mullet fry into the Arno river: Seasonal appearance, daily activity, and feeding rhythms. Fisheries Research, 1(??):219–234, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783681900266.

Unger:1989:SDC

[UB89a] Philip A. Unger and Stephen B. Brandt. Seasonal and diel changes in sampling conditions for acoustic surveys of fish abundance in small lakes. Fisheries Research, 7(4):353–366, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783689900672.

Uthe:1989:BRT

[UB89b] J. F. Uthe and J. M. Bewers. Book review: Toxic contamination in large lakes: Norbert W. Schmidtke (Editor), Lewis Publishers Inc., 4 vol. set, £158.60, Vol. 1, 364 pp., ISBN 0-87371-089-4; Vol. 2, 330 pp., ISBN 0-87371-090-8; Vol. 3, 440 pp., ISBN 0-87371-091-6; Vol. 4, 321 pp., ISBN 0-87371-092-4. Fisheries Research, 8 (2):200–202, November 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900337.

Ulltang:1987:PGI

[Ull87] Øyvind Ulltang. Potential gains from improved management of the Northeast Arctic cod stock. Fisheries Research, 5(2–3):319– 330, July 1987. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783687900506.

Vanderpuye:1981:FOD

[Van82] C. J. Vanderpuye. Further observation of the distribution and abundance of fish stocks in Volta Lake, Ghana. Fisheries Research, 1 (??):319–343, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900394.

Vondracek:1989:SSD

[VBBM89] Bruce Vondracek, Donald M. Baltz, Larry R. Brown, and Peter B. Moyle. Spatial, seasonal and diel distribution of fishes in a California reservoir dominated by native fishes. Fisheries Research, 7 (1–2):31–53, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783689900052.

VanDolah:1987:ERT

[VWN87] Robert F. Van Dolah, Priscilla H. Wendt, and Nick Nicholson. Effects of a research trawl on a hard-bottom assemblage of sponges and corals. Fisheries Research, 5(1):39–54, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783687900142.

Walker:1989:BRC

[Wal89] M. Greer Walker. Book review: The cell and molecular biology of fish oogenesis: S. S. Guraya, Monographs in Development Biology, Vol. 18. Karger, Basel, New York, 1986, 223 pp., 68 figs., £68.70, ISBN 3-8055-4253-4. Fisheries Research, 7(4): 386–388, June 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900714.

Warburton:1989:CVS

[War89] Kevin Warburton. Catchability variation in short beam trawl hauls. Fisheries Research, 7(1–2):55–63, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783689900064.

Watson:1981:SFE

[Wat82] Dwight J. Watson. Subsistence fish exploitation and implications for management in the Baram River system, Sarawak, Malaysia. Fisheries Research, 1(??):299–310, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783681900370.

Walch:1981:HRA

[WB82] Leonard A. Walch and E. P. Bergersen. Home range and activity patterns of lake trout in central Colorado. Fisheries Research, 1 (??):311–318, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900382.

Ward:1987:DPM

[WD87] T. J. Ward and T. L. O. Davis. Diel periodicity of Metanephrops australiensis and other deep-water crustaceans of northwest Australia. Fisheries Research, 5(1):91–97, May 1987. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578368790018X.

Weihs:1985:BRF

[Wei85] D. Weihs. Book review: Fish locomotion: by R. W. Blake. Cambridge University Press, New York, London, 1983, xx + 208 pp., illus., $49.50, ISBN 0-521-24303-3. Fisheries Research, 3(??):377– 378, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900384.

Wassenberg:1989:ETS

[WH89] T. J. Wassenberg and B. J. Hill. The effect of trawling and subsequent handling on the survival rates of the by-catch of prawn trawlers in Moreton Bay, Australia. Fisheries Research, 7(1–2):99– 110, February 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900106.

Wankowski:1988:PAS

[WHW88] Jacek W. J. Wankowski, Edward Hyduke, and Stephen Williams. Population age structure and mortality rates of jackass morwong (Nemadactylus macropterus) in Eastern Bass Strait, Australia. Fisheries Research, 6(4):317–335, September 1988. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0165783688900021.

Widdows:1986:BRF

[Wid86] J. Widdows. Book review: Fundamentals of aquatic toxicology: Gary M. Rand and Sam R. Petrocelli (Editors), Hemisphere Publishing, Washington, 1985, 666 pp., US $69.95, ISBN 0-89116-302-6. Fisheries Research, 4(3–4):355–356, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0165783686900147.

Wisely:1984:BRF

[Wis84] Baughan Wisely. Book review: Fish Farming Handbook. Food, Bait, Tropicals and Goldfish: by E. Evan Brown and John B. Gratzek. Avi Publishing Company, Westport, Connecticut, 1980, 390 pp., 150 illustrations, U.S. $19.00, ISBN 0-87055-341-0. Fisheries Research, 2(4):322–324, September 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783684900389.

Wadie:1985:EDS

[WR85] W. F. Wadie and F. A. Abdel Razek. The effect of damming of the shrimp population in the South-Eastern part of the Mediterranean Sea. Fisheries Research, 3(??):323–335, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368590030X.

Wakeman:1981:LCS

[WW82] John M. Wakeman and Donald E. Wohlschlag. Least-cost swimming speeds and transportation costs in some pelagic estuarine fishes. Fisheries Research, 1(??):117–127, ???? 1981– 1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368190014X.

Wyman:1981:BRM

[Wym82] Richard L. Wyman. Book review: Meeting on hydroacoustical methods for the estimation of marine fish populations: 25–29 June 1979. J. B. Suomala and K. I. Yudanov (Editors). Publ. Charles Stark Draper Laboratories, Cambridge, MA, U.S.A., Vol. 1, xi + 71 pp. and Vol. 2, xii + 964 pp. Fisheries Research, 1(??): 179–182, ???? 1981–1982. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783681900205.

Wyman:1983:BRE

[Wym83] R. L. Wyman. Book review: Ecology and ethology of fishes: Proceedings of the 2nd biennial symposium on the ethology and behavioral ecology of fishes, held at Normal, IL, U.S.A., October 19–22, 1979. Edited by D. L. G. Noakes and J. A. Ward. Printed as a special edition of Environmental Biology of Fishes 6 (1), 1981. Junk, The Hague, 144 pp., Dfl. 65.00/U.S.$34.00, ISBN 90-6193-821-X. Fisheries Research, 2(1):63–66, June 1983. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783683901042.

Wyman:1984:BRE

[Wym84] Richard L. Wyman. Book review: Echolocation in whales and dolphins: By P. E. Purves and G. E. Pilleri. Academic Press, London, 1983, 261 pp., 133 figs., price £26.60/$44.00, ISBN 0.12.567960.2. Fisheries Research, 2(3):219–222, April 1984. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783684900055.

Wyman:1986:BRE

[Wym86] Richard L. Wyman. Book review: Ecology of Barnegat Bay, New Jersey: M. J. Kennish and R. A. Lutz (Editors). Lecture Notes on Coastal and Estuarine Studies. Vol. 6. Springer Verlag, Berlin, Heidelberg, New York, 1984, 396 pp., soft cover DM 96.00, approx. US $37.70, ISBN 3-540-90935-4. Fisheries Research, 4(3– 4):358–359, December 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0165783686900160.

Yamane:1989:ECE

[YF89] Takeshi Yamane and Efren E. C. Flores. Evaluation of the catching efficiency of small pots for prawns. Fisheries Research, 8(1):

81–91, July 1989. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783689900428.

Young:1986:BRY

[You86] P. C. Young. Book review: From year to year. Interannual variability of the environment and fisheries of the Gulf of Alaska and the eastern Bering Sea: Warren S. Wooster (Editor). Washington Sea Grant, Seattle, WA, U.S.A., 1983, 208 pp., price US $8.00. Fisheries Research, 4(1):78–80, March 1986. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783686900317.

Zalewski:1985:EFD

[Zal85] Maciej Zalewski. The estimate of fish density and biomass in rivers on the basis of relationships between specimen size and efficiency of electrofishing. Fisheries Research, 3(??):147– 155, ???? 1985. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783685900153.

Zheng:1988:PDS

[ZW88] Jie Zheng and Carl J. Walters. Population dynamics and stock assessment of Wanshan Spring Decapterus maruadsi (T. & S.) in South China Sea. Fisheries Research, 6(3):217–231, May 1988. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016578368890015X.